### v2

# Sequential Eigenvalue Statistics for Change-Point Detection in Covariance Matrices

Abstract
Testing for change points in sequences of covariance matrices is an important and equally challenging problem in statistical methodology with applications in various fields. Motivated by the observation that even in cases where the ratio between dimension and sample size is as small as $0.05$, tests based on a fixed-dimension asymptotics do not keep their preassigned level, we propose to derive critical values of test statistics using an asymptotic regime where the dimension diverges at the same rate as the sample size. This paper introduces a novel and well-founded statistical methodology for detecting change points in a sequence of moderately dimensional covariance matrices. Our approach utilizes a min-type statistic based on a sequential process of likelihood ratio statistics. This is used to construct a test for the hypothesis of the existence of a change point with a corresponding estimator for its location. We provide theoretical guarantees by thoroughly analyzing the asymptotic properties of the sequential process of likelihood ratio statistics. In particular, we prove weak convergence towards a Gaussian process under the null hypothesis of no change. To identify the challenging dependency structure between consecutive test statistics, we employ tools from random matrix theory and stochastic processes.
sequential processes,
keywords:
[…]
1 Introduction
Having its origins in quality control (see Wald, 1945; Page, 1954, for two early references), change point detection has been an extremely active field of research until today with numerous applications in finance, genetics, seismology or sports to name just a few. In the last decade, a large part of the literature on change point detection considers the problem of detecting a change point in a high-dimensional sequence of means (see Jirak, 2015; Cho and Fryzlewicz, 2015; Dette and Gösmann, 2020; Enikeeva and Harchaoui, 2019; Liu et al., 2020, 2021; Chen et al., 2022; Wang et al., 2022; Zhang et al., 2022a, among many others).
Compared to the vast body of work on the change-point problem for a sequence of high-dimensional means, the literature on the problem of detecting structural breaks in the corresponding covariance matrices is relatively scarce. For the low dimensional setting we refer to Chen and Gupta (2004), Lavielle and Teyssiere (2006), Galeano and Peña (2007), Aue et al. (2009) and Dette and Wied (2016), among others, who study different methods and aspects of the change point problem under the assumption that the sample size converges to infinity while the dimension is fixed. We also refer to Theorem 1.1.2 in Csörgő and Horváth (1997) who provide a test statistic and its asymptotic distribution under the null hypothesis for normally distributed data. However, even in cases where the ratio between dimension and sample size is rather small, it can be observed that statistical guarantees derived from fixed-dimension asymptotics can be misleading. For instance, we display in Table 1 the simulated type I error of two commonly used tests for a change point in a sequence of covariance matrices. The first method (CH) is based on sequential likelihood ratio statistics, where the critical values have been determined by classical asymptotic arguments assuming that the dimension is fixed (see Theorem 1.1.2 in Csörgő and Horváth, 1997). The second approach (AHHR) is a test proposed by Aue et al. (2009), which is based on a quadratic form of the vectorized CUSUM statistic of the empirical covariance matrix. Again, the determination of critical values relies on fixed-dimensional asymptotics. We observe that even in the case where the ratio between the dimension and sample size is as small as $0.05$, the nominal level $\alpha=0.05$ of the CH test is exceeded by more than a factor of three. On the other hand, the AHHR test provides only a reasonable approximation of the nominal level if the ratio between dimension and sample size is $0.025$. Note that this test requires the inversion of an estimate of a large dimensional covariance matrix and is only applicable if the sample size is larger than the squared dimension.
Dimension 5 10 15 20 25
Empirical level CH 0.05 0.16 0.39 0.82 1.00
AHHR 0.03 0.01 0.00 - -
Table 1: Simulated type I errors of the sequential likelihood ratio test (Theorem 1.1.2 in Csörgő and Horváth, 1997) and the test of Aue et al. (2009) for a sample size of $n=200$ (500500 simulation runs, nominal level $\alpha=0.05$, standard normally distributed data). Critical values are determined by fixed dimension asymptotics. If ”-” is reported, the corresponding test is not applicable.
Meanwhile, several authors have also discussed the problem of estimating a change point in a sequence of covariance matrices in the high-dimensional regime. For example, Avanesov and Buzun (2018) propose a multiscale approach to estimate multiple change points, while Wang et al. (2021) investigate the optimality of binary and wild binary segmentation for multiple change point detection. We further mention the work of Dette et al. (2022), who propose a two-stage approach to detect the location of a change point in a sequence of very high-dimensional covariance matrices. Li and Gao (2024) pursue a similar approach to develop a change-point test for high-dimensional correlation matrices.
The literature on testing for change points is relatively scarce. In principle, one can develop change point analysis based on a vectorization of the covariance matrices using inference tools for a sequence of means. This approach essentially boils down to comparing the matrices before and after the change point with respect to a vector norm. However, in general, this approach does not yield an asymptotically distribution free test statistic. Moreover, as pointed out by Ryan and Killick (2023), such distances do not reflect the geometry induced on the space of positive definite matrices. Their work introduces a change-point test based on an alternative distance defined on the space of positive definite matrices, which compares sequentially the multivariate ratio $\boldsymbol{\Sigma}^{-1}_{1}\boldsymbol{\Sigma}_{2}$ of the two covariance matrices $\boldsymbol{\Sigma}_{1}$ and $\boldsymbol{\Sigma}_{2}$ before and after a potential change point with the identity matrix. As a consequence, under the null hypothesis of no change point, their test statistic is independent of the underlying covariance structure, which makes it possible to derive quantiles for statistical testing in the regime where the dimension diverges at the same rate as the sample size. However, the approach of these authors is based on a combination of a point-wise limit theorem from random matrix theory with a Bonferroni correction. Therefore, as pointed out in Section 4 of Ryan and Killick (2023), the resulting test may be conservative in applications. Moreover, this methodology is tailored to centered data, and it is demonstrated in Zheng et al. (2015), that an empirical centering introduces a non-negligible bias in the central limit theorem for the corresponding linear spectral statistic.
In this paper, we propose an alternative test for detecting a change point in a sequence of covariance matrices, which takes the strong dependence between consecutive test statistics into account to avoid the drawbacks of previous works. Our approach is based on a sequential process of likelihood ratio test (LRT) statistics, where the dimension of the data grows at the same rate as the sample size. We combine tools from random matrix theory and stochastic processes to develop and analyze statistical methodology for change point analysis in the covariance structure. Random matrix theory is a common tool to investigate asymptotic properties of LRT in moderately high-dimensional scenarios for classical testing problems. An early reference in this direction is Bai et al. (2009), who study one- and two-sample problems for covariance matrices and provide Gaussian approximations for LRTs in high dimensions. Moreover, Jiang and Yang (2013) establish central limit theorems for several classical LRT statistics under the null hypotheses. Both works rely on the normal assumption.
Since these seminal works, numerous researchers have investigated related problems (see Jiang and Qi, 2015; Dette and Dörnemann, 2020; Bao et al., 2022; Dörnemann, 2023; Heiny and Parolya, 2024, among others). None of these papers considers sequential LRT statistics to develop change point analysis. Moreover, our approach is conceptually different from most existing work on testing for change points in high-dimensional data (see, for example, Liu et al. (2021) for changes in a mean vector and Wang and Yao (2021) for changes in a covariance matrix) and does neither require a sparsity nor a sub-Gaussian assumption. Having this line of literature in mind, we can summarize the main contributions of this paper.
  • We propose a novel methodology to test for a change point in a sequence of moderately high-dimensional covariance matrices based on a minimum of sequential LRT statistics. Under the null hypothesis, this statistic admits a simple limiting distribution in the regime where the dimension diverges proportionally to the sample size. Unlike most other approaches, the distribution of the test statistic under the null hypothesis is invariant to the population covariance matrix. This result facilitates the introduction of a simple asymptotic testing procedure with favorable finite-sample properties. Most notably, our approach takes the strong dependence structure between consecutive test statistics into account, whose analysis has been recognized as a challenging problem in the literature (see Ryan and Killick, 2023), and which has not been addressed in previous works.
  • Investigating sequential statistics introduces new mathematical challenges compared to the analysis of the standard (non-sequential) LRT, namely (i) the convergence of the finite-dimensional distributions and (ii) the asymptotic tightness of the sequential log-LRT statistics. Indeed, the weak convergence result implied by (i) and (ii) is a novel, technically challenging contribution, given that sequential LRT statistics have not been studied in such a framework before.
    To establish (i), we derive an asymptotic representation of the test statistics and apply a martingale CLT to the dominating term in this decomposition. Note that for given time points $t_{1},t_{2}\in[0,1]$, the corresponding LRT statistics are highly correlated, and a nuanced analysis is required to determine their covariance. Regarding (ii), we show asymptotic equicontinuity of the sequential log-LRT statistics by deriving uniform inequalities for the moments of the increments of the process.
  • Along the way, we develop a consistent estimator of the kurtosis. As numerous results in random matrix theory and high-dimensional statistics demonstrate that spectral statistics depend critically on whether the kurtosis equals three (Bai and Silverstein, 2010; Zhang et al., 2022b; Pan and Zhou, 2008; Zheng, 2012; Yin et al., 2023), this estimator is believed to be of independent methodological interest.
The remaining part of this work is structured as follows. In Section 2, we present the new method to detect a change-point in a covariance structure of moderate dimension, and provide the main theoretical guarantees. In numerical experiments given in Section 3, we compare the finite-sample size properties of our test as well as the change-point estimator to other approaches. The proofs of our theoretical results are deferred to Section 4 and the supplementary material.
2 Change point analysis by a sequential LRT process
Let $\mathbf{y}_{1},\ldots,\mathbf{y}_{n}$ be a sample of independent random vectors such that $\mathbf{y}_{i}=(y_{1i},\ldots,y_{pi})^{\top}=\boldsymbol{\Sigma}_{i}^{1/2}\mathbf{x}_{i}$ for i.i.d. $p$-dimensional random vectors $\mathbf{x}_{i}$ and covariance matrices $\boldsymbol{\Sigma}_{i}=\boldsymbol{\Sigma}_{i,n}$, $1\leq i\leq n.$ We are interested in testing for a change in the covariance structure of $\mathbf{y}_{1},\ldots,\mathbf{y}_{n}$, and consider the hypotheses
$H_{0}:$ $\boldsymbol{\Sigma}_{1}=\ldots=\boldsymbol{\Sigma}_{n}$ (2.1)
versus
$H_{1}:$ $\boldsymbol{\Sigma}_{1}=\ldots=\boldsymbol{\Sigma}_{\lfloor nt^{\star}\rfloor}\neq\boldsymbol{\Sigma}_{\lfloor nt^{\star}\rfloor+1}=\ldots=\boldsymbol{\Sigma}_{n},$ (2.2)
where the location $t^{\star}\in(t_{0},1-t_{0})$ of the change point is unknown and $t_{0}>0$ is a positive constant. We define
$\hat{\boldsymbol{\Sigma}}_{i:j}^{\textnormal{cen}}$ $=\frac{1}{j-i}\sum\limits_{k=i}^{j}\left(\mathbf{y}_{k}-\overline{\mathbf{y}}_{i:j}\right)\left(\mathbf{y}_{k}-\overline{\mathbf{y}}_{i:j}\right)^{\top},\quad 1\leq i\leq j\leq n,$ (2.3)
as the sample covariance matrices calculated from the data $\mathbf{y}_{i},\ldots,\mathbf{y}_{j}$, where
$\overline{\mathbf{y}}_{i:j}=\frac{1}{j-i+1}\sum_{k=i}^{j}\mathbf{y}_{k}$
denotes the sample mean of $\mathbf{y}_{i},\ldots,\mathbf{y}_{j}$. Finally, we define
$\hat{\boldsymbol{\Sigma}}^{\textnormal{cen}}$ $=\hat{\boldsymbol{\Sigma}}_{1:n}^{\textnormal{cen}},$ (2.4)
as the sample covariance matrix calculated from the full sample and consider the statistic
$\Lambda_{n,t}^{\textnormal{cen}}=\frac{\big|\hat{\boldsymbol{\Sigma}}^{\textnormal{cen}}_{1:\lfloor nt\rfloor}\big|^{\frac{1}{2}\lfloor nt\rfloor}\big|\hat{\boldsymbol{\Sigma}}^{\textnormal{cen}}_{(\lfloor nt\rfloor+1):n}\big|^{\frac{1}{2}(n-\lfloor nt\rfloor)}}{\big|\hat{\boldsymbol{\Sigma}}^{\textnormal{cen}}\big|^{\frac{1}{2}n}},\quad t\in(0,1).$ (2.5)
[…]
To formulate the statistical properties of this process, we make the following assumptions.
  1. (A-1)
    $y_{n}=p/n\to y\in(0,1)$ as $n\to\infty$ such that $y<t_{0}\wedge(1-t_{0})$ for some $t_{0}\in(0,1)$.
  2. (A-2)
    The components $x_{ji}$ of the vector $\mathbf{x}_{i}$ are i.i.d. with respect to some continuous distribution ($1\leq i\leq n,1\leq j\leq p$), and satisfy $\mathbb{E}[x_{11}^{2}]=1$, $\mathbb{E}[x_{11}^{4}]>1$ and $\mathbb{E}|x_{11}|^{4+\delta}<\infty$ for some $\delta>0$.
  3. (A-3)
    We have uniformly with respect to $n\in\mathbb{N}$
    $0<\lambda_{\min}(\boldsymbol{\Sigma}_{1})\leq\lambda_{\max}(\boldsymbol{\Sigma}_{1})<\infty.$
An important ingredient for an appropriate centering of $\log\Lambda_{n,t}^{\textnormal{cen}},$ is an estimator of the kurtosis
$\kappa_{n}=\mathbb{E}[x_{11}^{4}]$
of the unobserved random variable $x_{11}$, which can be represented by formula (9.8.6) in Bai and Silverstein (2010) in the form
$\kappa_{n}=3+\frac{\operatorname{Var}\left(\|\mathbf{y}_{1}-\mathbb{E}[\mathbf{y}_{1}]\|_{2}^{2}\right)-2\|\boldsymbol{\Sigma}\|_{F}^{2}}{\sum_{j=1}^{p}\Sigma_{jj}^{4}}.$ (2.6)
For its estimation, we therefore introduce the quantities
$\hat{\tau}_{n}$ $=\operatorname{tr}\big((\hat{\boldsymbol{\Sigma}}^{\textnormal{cen}})^{2}\big)-\frac{1}{n}\big(\operatorname{tr}\hat{\boldsymbol{\Sigma}}^{\textnormal{cen}}\big)^{2},$
$\hat{\nu}_{n}$ $=\frac{1}{n-1}\sum_{i=1}^{n}\left(\|\mathbf{y}_{i}-\overline{\mathbf{y}}\|_{2}^{2}-\frac{1}{n}\sum_{i^{\prime}=1}^{n}\|\mathbf{y}_{i^{\prime}}-\overline{\mathbf{y}}\|_{2}^{2}\right)^{2},$
$\hat{\omega}_{n}$ $=\sum_{j=1}^{p}\Big\{\frac{1}{n}\sum_{i=1}^{n}\Big(y_{ji}-\frac{1}{n}\sum_{i^{\prime}=1}^{n}y_{ji^{\prime}}\Big)^{2}\Big\}^{2}.$
and define the estimator
$\hat{\kappa}_{n}$ $=\max\Big\{3+\frac{\hat{\nu}_{n}-2\hat{\tau}_{n}}{\hat{\omega}_{n}},1\Big\}.$
Our first result provides the consistency of $\hat{\kappa}_{n}$ for $\kappa_{n}$ under the null hypothesis. Its proof is postponed to Section A.4.
Proposition 1.
Suppose that assumptions (A-2)-(A-3) are satisfied, and $p/n\to y\in(0,\infty)$ as $n\to\infty$. Then, under $H_{0}$, we have
$\frac{\hat{\kappa}_{n}}{\kappa_{n}}\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}1.$
Remark 1.
In the case $\mathbb{E}[y_{11}]=0$, a related estimator for $\kappa_{n}$ was proposed by Lopes et al. (2019). To the best of our knowledge, $\hat{\kappa}_{n}$ is the first estimator to be equipped with theoretical guarantees under general (possibly nonzero) means in the regime where the dimension is asymptotically proportionally to the sample size. This estimator is believed to be of independent methodological interest. Indeed, the excess kurtosis $\kappa_{n}-3$ is a key quantity in extending asymptotic results for spectral statistics from the Gaussian case to non-Gaussian settings, see, e.g., Bai and Silverstein (2010); Pan and Zhou (2008); Zheng (2012); Yin et al. (2023); Najim and Yao (2016); Zhang et al. (2022b). Since this parameter directly affects the limiting behavior of eigenvalue-based statistics, we expect that $\hat{\kappa}_{n}$ will find applications in testing problems for moderately high-dimensional data.
We will show that under the null hypothesis we can approximate the expected value and the variance of $2\log\Lambda_{n,t}^{\textnormal{cen}}$ by
$\tilde{\mu}_{n,t}$ $=n\Big(n-p-\frac{3}{2}\Big)\log\Big(1-\frac{p}{n-1}\Big)-\lfloor nt\rfloor\Big(\lfloor nt\rfloor-p-\frac{3}{2}\Big)\log\Big(1-\frac{p}{\lfloor nt\rfloor-1}\Big)$
$\quad-(n-\lfloor nt\rfloor)\Big(n-\lfloor nt\rfloor-p-\frac{3}{2}\Big)\log\Big(1-\frac{p}{n-\lfloor nt\rfloor-1}\Big)+\frac{(\hat{\kappa}_{n}-3)p}{2}$ (2.7)
and
$\sigma_{n,t}^{2}$ $=2\log\left(1-\frac{p}{n}\right)-2\left(\frac{\lfloor nt\rfloor}{n}\right)^{2}\log\left(1-\frac{p}{\lfloor nt\rfloor}\right)-2\left(\frac{n-\lfloor nt\rfloor}{n}\right)^{2}\log\left(1-\frac{p}{n-\lfloor nt\rfloor}\right)$ (2.8)
respectively. With these quantities we consider the standardized LRT and define the min-type statistic
$M_{n}^{\textnormal{cen}}=\min_{t\in[t_{0},1-t_{0}]}\frac{2\log\Lambda_{n,t}^{\textnormal{cen}}-\tilde{\mu}_{n,t}}{n\sigma_{n,t}}~.$
In the next theorem, we provide the limiting distribution of $M_{n}^{\textnormal{cen}}$ under the null hypothesis of no change.
Theorem 1.
If Assumption (A-1), (A-2) with $\delta>4$ and Assumption (A-3) are satisfied, then we have under $H_{0}$
$M_{n}^{\textnormal{cen}}\stackrel{{\scriptstyle\mathcal{D}}}{{\to}}\min_{t\in[t_{0},1-t_{0}]}\frac{Z(t)}{\sqrt{\sigma(t,t)}},$ (2.9)
where $(Z(t))_{t\in[t_{0},1-t_{0}]}$ denotes a centered Gaussian process with covariance kernel
$\sigma(t_{1},t_{2})$ $=\operatorname{cov}(Z(t_{1}),Z(t_{2}))$
$=2\log(1-y)-2t_{1}t_{2}\log(1-y/t_{2})-2(1-t_{1})t_{2}\log\Big(1-\frac{(t_{2}-t_{1})y}{(1-t_{1})t_{2}}\Big)$ (2.10)
$\quad-2(1-t_{1})(1-t_{2})\log(1-y/(1-t_{1}))$
for $t_{0}\leq t_{1}\leq t_{2}\leq 1-t_{0}.$
A proof of this result can be found in Section 4.1. Note that the limiting distribution in (2.9) contains no nuisance parameters. Consequently, if $q_{\alpha}$ denotes the $\alpha$-quantile of the limit distribution, the decision rule, which rejects the null hypothesis in (2.1), whenever
$M_{n}^{\textnormal{cen}}<q_{\alpha}.$ (2.11)
defines an asymptotic level $\alpha$-test for the hypotheses of a change point in the sequence $\boldsymbol{\Sigma}_{1},\ldots,\boldsymbol{\Sigma}_{n}$. The quantile $q_{\alpha}$ can be found numerically, replacing the asymptotic ratio $y$ in (2.10) by $p/n$. Then, Theorem 1 implies that the level of the test (2.11) can be asymptotically controlled under $H_{0}$, that is,
$\lim_{n\to\infty}\mathbb{P}\left(M_{n}^{\textnormal{cen}}<q_{\alpha}\right)=\alpha.$
3 Finite-sample properties
The necessity of $t_{0}$. The parameter $t_{0}$ ensures the applicability of the likelihood-ratio principle and is determined by the user. Parameters of this type appear frequently in monitoring high-dimensional covariance structures (see, for example, Ryan and Killick, 2023; Dörnemann and Dette, 2024; Dörnemann and Paul, 2024). In fact, there is one-to-one correspondence between $t_{0}$ and the minimum segment length parameter $\ell$ in Ryan and Killick (2023), and thus $t_{0}$ underlies the same paradigm as $\ell$ outlined in the aforementioned work. On the one hand, small values of $t_{0}$ are likely to increase the type-I error. In such cases, the maximal statistic will be dominated by covariance estimates corresponding to potential change points $t$ close to $p/n$ (or, by symmetry, close to 1p/n1-p/n) which admit large eigenvalues. On the other hand, in many applications, the user may want to avoid large values for $t_{0}$, as such choices shrink the localization interval for change-point candidates. Therefore, it is important to understand how small the tuning parameter $t_{0}$ can be chosen without affecting the performance of the proposed method. Regarding the selection of $t_{0}$, it should first be noted that the parameter is unitless and does not need to be adapted to the scale of the model. By the design of the test statistic, a necessary lower bound will be $t_{0}>p/n\vee(1-p/n).$ In our simulation study, we found that the testing method is stable if $t_{0}>(p/n+0.05)\vee 0.2$. If the user is primarily interested in estimating the change point location, they may select $t_{0}$ closer to the critical threshold $p/n\vee(1-p/n)$.
Estimating the change-point location. If $H_{0}$ is rejected by the test (2.11), it is natural to ask for the location of the change point. For this purpose, we propose the following estimator:
$\hat{\tau}^{\star}\in\operatorname{argmin}_{t\in[t_{0},1-t_{0}]}\frac{2\log\Lambda^{\textnormal{cen}}_{n,t}-\tilde{\mu}_{n,t}}{n}.$ (3.1)
In Section 3.2, we investigate the numerical performance of $\hat{\tau}^{\star}$ and compare it to the estimators of Aue et al. (2009) and Ryan and Killick (2023).
3.1 Numerical experiments for change-point detection
In the following, we provide numerical results on the performance of the new test (2.11) in comparison to the test proposed by Ryan and Killick (2023). All reported results are based on $500$ simulation runs, and the nominal level is $\alpha=0.05$. The change-point location is chosen as $t^{\star}=0.5$.
Recall that we observe the data $\mathbf{y}_{i}=\boldsymbol{\Sigma}_{i}^{1/2}\mathbf{x}_{i}$ for $1\leq i\leq n$, where $\boldsymbol{\Sigma}_{i}^{1/2}$ and $\mathbf{x}_{i}$ are not directly observed. We first consider independent standard normal distributed entries ($x_{11}\sim\mathcal{N}(0,1)$) in the vectors $\mathbf{x}_{i}$ and
$\boldsymbol{\Sigma}_{1}=\mathbf{I},\quad\boldsymbol{\Sigma}_{n}=\operatorname{diag}(1,\ldots,1,\underbrace{\eta,\ldots,\eta}_{p/2}),\quad\eta\geq 1,$~ (3.2)
as the covariance matrices before and after the change point, where the case $\eta=1$ corresponds to null hypothesis (2.1). The empirical rejection probabilities of the test (2.11) are displayed in the left panels of Figure 1 for $(n,p)=(600,50)$ (first row) and $(600,80)$ (middle row) and $(800,100)$ (third row) and various values of $\eta$. We observe that the test keeps its nominal level well and that the power increases quickly with $\eta$. For the sake of comparison, we also display the empirical rejection probabilities of the test proposed in Ryan and Killick (2023). As stated by these authors, this test is conservative, and we observe a substantial improvement with respect to power by the new test (2.11), which takes the dependencies of the statistics $\Lambda_{n,t}^{\textnormal{cen}}$ for different values of $t$ into account.
Next, we consider an adaptation of (3.3), where the matrix $\boldsymbol{\Sigma}_{n}$ is randomly generated with a prescribed spectrum, that is
$\boldsymbol{\Sigma}_{1}=\mathbf{I},\quad\boldsymbol{\Sigma}_{n}=\mathbf{U}_{\eta}\operatorname{diag}(1,\ldots,1,\underbrace{\eta,\ldots,\eta}_{p/2})\mathbf{U}_{\eta}^{\top},\quad\eta\geq 1,$ (3.3)
where $\mathbf{U}_{0}=\mathbf{I}$, and $\mathbf{U}_{\eta}$ are independent random matrices uniformly distributed on the orthogonal group for $\eta>1$. The independent entries in the matrix $\mathbf{X}$ are generated from a (uniform) $\mathcal{U}(0,1)$-distribution. The corresponding results are displayed in the right panels of Figure 1. Comparing these results with the left panels, we observe that the approximation of the nominal levels in the two models (3.2) and (3.3) is comparable. The new test shows a favorable performance under both alternatives (3.2) and (3.3) Notably, we observe an increase in power for (3.3) compared to (3.2), as (3.3) involves changes in both eigenvalues and eigenvectors, whereas (3.2) has changes only in the eigenvalues. In all cases under consideration, the new test outperforms the conservative method proposed by Ryan and Killick (2023) in terms of level approximation and power increase.
Figure 1: Empirical rejection rates of the new test (2.11) (triangle) compared to the test of Ryan and Killick (2023) (diamond), where $t_{0}=0.2,~t^{\star}=0.5$ and $(n,p)=(600,50)$ (first row), $(n,p)=(600,80)$ (middle row), $(n,p)=(800,100)$ (third row). Left panels: model (3.2)), where $x_{11}\sim{\cal N}(0,1)$. Right panels: model (3.3), where $x_{11}\sim\mathcal{U}(0,1)$.
Next, we investigate the performance of our method in the case where the change only affects the eigenvectors, but not the eigenvalues of $\boldsymbol{\Sigma}_{1}$ and $\boldsymbol{\Sigma}_{n}$. We consider the covariance matrices
$\boldsymbol{\Sigma}_{1}=\mathbf{Q}_{1}\operatorname{diag}(\underbrace{2,\ldots,2}_{p/2},\underbrace{1,\ldots,1}_{p/2})\mathbf{Q}_{1}^{\top},\quad\boldsymbol{\Sigma}_{n}=\mathbf{Q}_{2}\operatorname{diag}(\underbrace{2,\ldots,2}_{p/2},\underbrace{1,\ldots,1}_{p/2})\mathbf{Q}_{2}^{\top},$ (3.4)
where $\mathbf{Q}_{1},\mathbf{Q}_{2}$ are independent random matrices uniformly distributed on the orthogonal group. Note that this scenario corresponds to the alternative $H_{1}$ with overwhelming probability. Moreover, $\boldsymbol{\Sigma}_{1}$ and $\boldsymbol{\Sigma}_{n}$ share the same eigenvalues, so the spectrum does not change. Under $H_{0},$ we set $\boldsymbol{\Sigma}_{1}=\boldsymbol{\Sigma}_{n}$, and $\boldsymbol{\Sigma}_{1}$ is generated as above in (3.4). In Table 2, we display the empirical rejection rates of our proposed test. We observe that the proposed test attains full power against such alternatives.
$(n,p)$ (600,80) (800,100)
$x_{11}\sim\mathcal{U}(0,1)$ $H_{0}$ 0.058 0.066
$H_{1}$ 1.000 1.000
$x_{11}\sim\mathcal{N}(0,1)$ $H_{0}$ 0.068 0.062
$H_{1}$ 1.000 1.000
Table 2: Empirical rejection rates of the new test (2.11) for different values of $(n,p)$ and distributions for $x_{11}$, where $t_{0}=0.2,~t^{\star}=0.5$
$\eta$ 1.1 1.2 1.3 1.4 1.5 1.6 1.7 1.8 1.9 2
$\hat{\tau}^{\star}$ mean 0.499 0.499 0.506 0.493 0.496 0.494 0.497 0.496 0.499 0.499
sd 0.123 0.112 0.091 0.066 0.054 0.034 0.025 0.018 0.010 0.008
MSE 0.015 0.012 0.008 0.004 0.003 0.001 0.001 0.000 0.000 0.000
RK mean 0.529 0.497 0.425 0.422 0.436 0.462 0.473 0.479 0.487 0.491
sd 0.203 0.198 0.157 0.105 0.073 0.048 0.042 0.033 0.023 0.016
MSE 0.042 0.039 0.030 0.017 0.009 0.004 0.002 0.001 0.001 0.000
$\hat{\tau}^{\star}$ mean 0.507 0.499 0.497 0.496 0.498 0.496 0.497 0.498 0.497 0.498
sd 0.129 0.111 0.097 0.073 0.055 0.039 0.024 0.016 0.011 0.008
MSE 0.017 0.012 0.009 0.005 0.003 0.002 0.001 0.000 0.000 0.000
RK mean 0.515 0.501 0.420 0.395 0.413 0.431 0.450 0.460 0.468 0.474
sd 0.211 0.210 0.161 0.096 0.074 0.058 0.044 0.040 0.033 0.026
MSE 0.045 0.044 0.032 0.020 0.013 0.008 0.004 0.003 0.002 0.001
$\hat{\tau}^{\star}$ mean 0.495 0.495 0.494 0.494 0.496 0.498 0.499 0.499 0.499 0.500
sd 0.124 0.106 0.082 0.055 0.038 0.023 0.013 0.010 0.006 0.005
MSE 0.015 0.011 0.007 0.003 0.001 0.001 0.000 0.000 0.000 0.000
RK mean 0.551 0.486 0.414 0.404 0.429 0.451 0.464 0.471 0.476 0.483
sd 0.204 0.199 0.136 0.077 0.057 0.038 0.033 0.027 0.024 0.019
MSE 0.044 0.040 0.026 0.015 0.008 0.004 0.002 0.002 0.001 0.001
Table 3: Simulated mean, standard deviation and mean squared error of the estimator $\hat{\tau}^{\star}$ in (3.1) and the estimator proposed in Ryan and Killick (2023) (RK), where $t_{0}=0.2,~t^{\star}=0.5$. The model is given by (3.2) with independent $\mathcal{N}(0,1)$-distributed entries in the matrix $\mathbf{X}$ and $(n,p)=(600,50)$ (top), $(n,p)=(600,80)$ (middle) and $(n,p)=(800,100)$ (bottom)
$\eta$ 1.1 1.2 1.3 1.4 1.5 1.6 1.7 1.8 1.9 2
$\hat{\tau}^{\star}$ mean 0.507 0.499 0.498 0.496 0.498 0.500 0.499 0.500 0.500 0.500
sd 0.123 0.087 0.055 0.033 0.009 0.006 0.004 0.003 0.003 0.001
MSE 0.015 0.008 0.003 0.001 0.000 0.000 0.000 0.000 0.000 0.000
RK mean 0.548 0.518 0.454 0.473 0.490 0.495 0.497 0.498 0.498 0.498
sd 0.195 0.193 0.106 0.048 0.021 0.012 0.007 0.006 0.006 0.007
MSE 0.040 0.037 0.013 0.003 0.001 0.000 0.000 0.000 0.000 0.000
$\hat{\tau}^{\star}$ mean 0.496 0.501 0.497 0.498 0.497 0.499 0.500 0.500 0.500 0.500
sd 0.118 0.091 0.056 0.033 0.018 0.008 0.004 0.004 0.002 0.001
MSE 0.014 0.008 0.003 0.001 0.000 0.000 0.000 0.000 0.000 0.000
RK mean 0.548 0.505 0.428 0.458 0.481 0.490 0.495 0.496 0.497 0.498
sd 0.198 0.199 0.109 0.052 0.032 0.019 0.010 0.008 0.008 0.005
MSE 0.041 0.039 0.017 0.004 0.001 0.000 0.000 0.000 0.000 0.000
$\hat{\tau}^{\star}$ mean 0.495 0.501 0.499 0.501 0.499 0.499 0.500 0.500 0.500 0.500
sd 0.106 0.075 0.036 0.015 0.007 0.006 0.001 0.002 0.001 0.001
MSE 0.011 0.006 0.001 0.000 0.000 0.000 0.000 0.000 0.000 0.000
RK mean 0.590 0.475 0.443 0.474 0.492 0.496 0.498 0.498 0.499 0.499
sd 0.187 0.176 0.069 0.038 0.016 0.009 0.006 0.004 0.003 0.003
MSE 0.043 0.032 0.008 0.002 0.000 0.000 0.000 0.000 0.000 0.000
Table 4: Simulated mean, standard deviation and mean squared error of the estimator $\hat{\tau}^{\star}$ in (3.1) and the estimator proposed in Ryan and Killick (2023) (RK), where $t_{0}=0.2,~t^{\star}=0.5$. The model is given by (3.3) with independent $\mathcal{U}(0,1)$-distributed entries in the matrix $\mathbf{X}$ and $(n,p)=(600,50)$ (top), $(n,p)=(600,80)$ (middle) and $(n,p)=(800,100)$ (bottom).
$\eta$ 1.1 1.2 1.3 1.4 1.5 1.6 1.7 1.8 1.9 2
$\hat{\tau}^{\star}$ mean 0.497 0.491 0.475 0.452 0.446 0.428 0.421 0.415 0.415 0.407
sd 0.167 0.152 0.150 0.133 0.124 0.111 0.081 0.080 0.066 0.048
MSE 0.037 0.031 0.028 0.020 0.017 0.013 0.007 0.007 0.005 0.002
RK mean 0.475 0.455 0.454 0.417 0.401 0.359 0.358 0.367 0.345 0.381
sd 0.287 0.278 0.272 0.247 0.231 0.201 0.176 0.155 0.128 0.114
MSE 0.088 0. 080 0.077 0.061 0.053 0.042 0.033 0.025 0.020 0.015
AHHR mean 0.512 0.516 0.507 0.512 0.506 0.495 0.486 0.486 0.482 0.478
sd 0.087 0.088 0.085 0.083 0.079 0.076 0.071 0.071 0.073 0.068
MSE 0.020 0.021 0.019 0.019 0.017 0.015 0.012 0.012 0.012 0.011
$\hat{\tau}^{\star}$ mean 0.507 0.482 0.468 0.452 0.447 0.432 0.420 0.419 0.414 0.410
sd 0.158 0.159 0.147 0.141 0.132 0.109 0.090 0.082 0.068 0.058
MSE 0.036 0.032 0.026 0.022 0.020 0.013 0.009 0.007 0.005 0.003
RK mean 0.483 0.463 0.451 0.409 0.407 0.379 0.365 0.358 0.354 0.355
sd 0.294 0.301 0.293 0.275 0.255 0.237 0.211 0.196 0.183 0.165
MSE 0.093 0.095 0.088 0.076 0.065 0.057 0.046 0.040 0.036 0.029
AHHR mean 0.505 0.507 0.507 0.503 0.502 0.499 0.504 0.495 0.499 0.491
sd 0.066 0.061 0.060 0.056 0.060 0.058 0.056 0.057 0.052 0.058
MSE 0.015 0.015 0.015 0.014 0.014 0.013 0.014 0.012 0.012 0.012
Table 5: Estimated change point location given by $\hat{\tau}^{\star}$ compared to Ryan and Killick (2023) (RK) and Aue et al. (2009) (AHHR) under model (3.2) based on 500 simulation runs in the setting $(n,p)=(200,10)$ (top), $(n,p)=(200,15)$ (bottom), $t_{0}=0.1,~t^{\star}=0.4,x_{11}\sim\mathcal{N}(0,1)$.
3.2 Numerical experiments for the change-point estimation
In this section, we compare the new change point estimator $\hat{\tau}^{\star}$ in (3.1) with the estimators proposed by Aue et al. (2009) (AHHR) and Ryan and Killick (2023) (RK). All results are again based on $500$ simulation runs.
In Table 3, we compare the mean, standard deviation and mean squared error of the new estimator $\hat{\tau}^{\star}$ in (3.1) with the RK estimator for the different alternatives in model (3.2) (with $\mathcal{N}(0,1)$-distributed independent entries in the matrix $\mathbf{X}$), where $t^{\star}=0.5$, $(n,p)=(600,50)$ (top), $(n,p)=(600,80)$ (middle) and $(n,p)=(800,100)$ (bottom). Note that the dimension is of comparable magnitude to the sample size, and therefore, the AHHR estimator cannot be computed and is therefore not included in the comparison. For example, for a dimension $p=50$, one requires at least a sample size of $(p+1)p/2+1=1276$ to calculate this estimator (some results for the AHHR estimator can be found in Table 5). We observe from the upper part of Table 3 that the new estimator (3.1) outperforms the RK estimator in all three cases under consideration. The smaller mean squared error of the new estimator (3.1) is caused by both a smaller bias and variance. In particular, the RK estimator admits a significant bias for moderately strong signals $\eta\approx 1.5.$
In Table 4, we display the results of the two estimators for model (3.3) with uniformly distributed data. The results are similar to those presented in Table 3 for model (3.2). Again, our method outperforms the alternative RK approach in terms of smaller mean squared error.
We conclude this section with a small comparison of the two estimators $\hat{\tau}^{\star}$ and RK with the estimator proposed by Aue et al. (2009) (AHHR) in the model (3.2). For this purpose, we select $t^{\star}=0.4$, and display the characteristics of the three change point estimators in Table 5. Note that the AHHR estimator can only be computed if the sample size is at least $p(p+1)/2+1$ and for this reason, we consider the cases $(n,p)=(200,10)$ (top), $(n,p)=(200,15)$ (bottom). As the dimension is relatively small compared to the sample size, we choose $t_{0}=0.1$. We observe that, even in such cases, the estimator AHHR admits a significant bias resulting in a larger MSE compared to the other two methods. Interestingly, the bias of RK increases as the signal strength $\eta$ increases from moderately to large values. In contrast, the new estimator $\hat{\tau}^{\star}$ has decreasing bias and standard deviation as $\eta$ increases. Moreover, the new estimator always outperforms RK indicated by a smaller mean squared error, and AHHR in the case $(n,p)=(200,15).$ For $(n,p)=(200,10),$ we observe that the mean squared error of AHHR is smaller for weak signal strength $\eta.$ However, even for large $\eta$, this method admits a significant bias and is therefore outperformed by our method.
4 Proofs of main results under the null hypothesis
4.1 Proof of Theorem 1
Throughout this section, we may assume $\mathbb{E}[x_{11}]=0$ by definition of $\Lambda_{n,t}^{\textnormal{cen}}$ without loss of generality. The first step in the proof of Theorem 1 consists of reducing it to a corresponding statement for the non-centered sample covariance matrix. For this purpose, we proceed with some preparations and define the non-centered sequential sample covariance matrices as
$\hat{\boldsymbol{\Sigma}}_{i:j}^{(n)}$ $=\hat{\boldsymbol{\Sigma}}_{i:j}=\frac{1}{j-i+1}\sum\limits_{k=i}^{j}\mathbf{y}_{j}\mathbf{y}_{j}^{\top},\quad 1\leq i\leq j\leq n,$
$\hat{\boldsymbol{\Sigma}}$ $=\hat{\boldsymbol{\Sigma}}_{1:n}.$
Consider the sequential likelihood ratio statistics
$\Lambda_{n,t}=\frac{\big|\hat{\boldsymbol{\Sigma}}_{1:\lfloor nt\rfloor}\big|^{\frac{1}{2}\lfloor nt\rfloor}\big|\hat{\boldsymbol{\Sigma}}_{(\lfloor nt\rfloor+1):n}\big|^{\frac{1}{2}(n-\lfloor nt\rfloor)}}{\big|\hat{\boldsymbol{\Sigma}}\big|^{\frac{1}{2}n}},\quad t\in(0,1).$ (4.1)
and the corresponding centered process
$\boldsymbol{\Lambda}_{n}=((2\log\Lambda_{n,t}-\mu_{n,t})/n)_{t\in[t_{0},1-t_{0}]},$
where the centering term is defined as
$\mu_{n,t}$ $=n\left(n-p-\frac{1}{2}\right)\log\Big(1-\frac{p}{n}\Big)-\lfloor nt\rfloor\left(\lfloor nt\rfloor-p-\frac{1}{2}\right)\log\Big(1-\frac{p}{\lfloor nt\rfloor}\Big)$
$\quad-(n-\lfloor nt\rfloor)\left(n-\lfloor nt\rfloor-p-\frac{1}{2}\right)\log\left(1-\frac{p}{n-\lfloor nt\rfloor}\right)+\frac{(\hat{\kappa}_{n}-3)p}{2},\quad t\in[t_{0},1-t_{0}].$
In the following theorem, we provide the convergence of the finite-dimensional distributions of $(\boldsymbol{\Lambda}_{n})_{n\in\mathbb{N}}.$
Theorem 2.
Suppose that assumptions (A-1), (A-2) for some $\delta>0$, and (A-3) are satisfied, and that $\mathbb{E}[x_{11}]=0$. For $n\to\infty$ and all fixed $k\in\mathbb{N}$, $t_{1},\ldots,t_{k}\in[t_{0},1-t_{0}]$, we have under $H_{0}$
$\Big(\frac{2\log\Lambda_{n,t_{i}}-\mu_{n,t_{i}}}{n}\Big)_{1\leq i\leq k}\stackrel{{\scriptstyle\mathcal{D}}}{{\to}}(Z(t_{i}))_{1\leq i\leq k},$
where $(Z(t))_{t\in[t_{0},1-t_{0}]}$ denotes the Gaussian process defined in Theorem 1.
The asymptotic tightness of $(\boldsymbol{\Lambda}_{n})_{n\in\mathbb{N}}$ is given in the next theorem.
Theorem 3.
Suppose that Assumptions (A-1), (A-2) with $\delta>4$ and (A-3) are satisfied, and that $\mathbb{E}[x_{11}]=0$. Then, the sequence $(\boldsymbol{\Lambda}_{n})_{n\in\mathbb{N}}$ is asymptotically tight in the space $\ell^{\infty}([t_{0},1-t_{0}]).$
Proofs of these statements can be found in Section 4.1.1 and A.2, respectively. Then, the weak convergence of $(\boldsymbol{\Lambda}_{n})_{n\in\mathbb{N}}$ towards a Gaussian process follows from the convergence of the finite-dimensional distributions (Theorem 2) and the tightness result (Theorem 3).
Corollary 1.
Suppose that assumptions (A-1), (A-2) with $\delta>4$, (A-3) are satisfied, and that $\mathbb{E}[x_{11}]=0$. Then, we have under the null hypothesis $H_{0}$ of no change point
$\Big(\frac{2\log\Lambda_{n,t}-\mu_{n,t}}{n}\Big)_{t\in[t_{0},1-t_{0}]}\stackrel{{\scriptstyle\mathcal{D}}}{{\to}}\big(Z(t)\big)_{t\in[t_{0},1-t_{0}]}\quad\textnormal{in }\ell^{\infty}([t_{0},1-t_{0}]),$
where $(Z(t))_{t\in[t_{0},1-t_{0}]}$ denotes the centered Gaussian process defined in Theorem 1.
Before continuing with the proof of Theorem 1, we comment on the integration of our theoretical result in the existing line of literature.
Remark 2.
  1. (1)
    Theorem 1 and Corollary 1 continue the line of literature on substitution principles in random matrix theory. When considering the spectral statistics of $\hat{\boldsymbol{\Sigma}}$ and $\hat{\boldsymbol{\Sigma}}^{(\textnormal{cen})}$, it was found by Zheng et al. (2015) that their asymptotic distributions are linked by a substitution principle. This results says that one needs to substitute the location parameter $c_{n}$ in the CLT for the linear spectral statistics of $\hat{\boldsymbol{\Sigma}}$ by $c_{n-1}$ to account for the centralization in $\hat{\boldsymbol{\Sigma}}^{\textnormal{cen}}.$ A similar result has been found by Yin et al. (2023) for the linear eigenvalue statistics of the sample correlation matrix. However, it is important to emphasize that the test statistic $\Lambda_{n,t}^{\textnormal{cen}}$ considered in this work is a functional of several strongly dependent eigenvalue statistics, and therefore these results are not applicable. In fact, the analysis of $\Lambda_{n,t}^{\textnormal{cen}}$ requires a careful study, accounting for its intricate structure. These challenges will be faced even when restricting our focus to the case of one-dimensional distributions of $(\Lambda_{n,t})_{t}$, let alone considering the process convergence.
  2. (2)
    For the process convergence of $(\boldsymbol{\Lambda}_{n})_{n\in\mathbb{N}}$ in the space of bounded functions, the stronger moments condition (A-2) with $\delta>4$ is needed, whereas moments of order $4+\delta$ for some $\delta>0$ are sufficient for the convergence of the finite-dimensional distributions of $(\log\Lambda_{n,t})_{t\in[t_{0},1-t_{0}]}$.
With these preparations, we are in a position to prove Theorem 1.
Proof of Theorem 1.
Note that
$\hat{\boldsymbol{\Sigma}}^{\textnormal{cen}}=\frac{n}{n-1}\hat{\boldsymbol{\Sigma}}-\overline{\mathbf{y}}\overline{\mathbf{y}}^{\top},$
where $\overline{\mathbf{y}}=\overline{\mathbf{y}}_{1:n}$ denotes the sample mean of $\mathbf{y}_{1},\ldots,\mathbf{y}_{n}$. Using the matrix determinant lemma, this implies
$\log|\hat{\boldsymbol{\Sigma}}^{\textnormal{cen}}|$ $=\log\big|\frac{n}{n-1}\hat{\boldsymbol{\Sigma}}\big|+\log\left(1-\overline{\mathbf{y}}^{\top}\hat{\boldsymbol{\Sigma}}^{-1}\overline{\mathbf{y}}\right)$
$=-p\log\left(1-\frac{1}{n}\right)+\log\left|\hat{\boldsymbol{\Sigma}}\right|+\log\left(1-\overline{\mathbf{y}}^{\top}\hat{\boldsymbol{\Sigma}}^{-1}\overline{\mathbf{y}}\right).$
A Taylor expansion shows that $p\log\left(1-\frac{1}{n}\right)=p/n+o(1)$, and it also holds $\log\left(1-\overline{\mathbf{y}}^{\top}\hat{\boldsymbol{\Sigma}}^{-1}\overline{\mathbf{y}}\right)=\log(1-p/n)+o(1)$ almost surely (see Section 4.3.1 in Heiny and Parolya, 2024). Thus, we obtain
$\log|\hat{\boldsymbol{\Sigma}}^{\textnormal{cen}}|$ $=\log\left|\hat{\boldsymbol{\Sigma}}\right|+\frac{p}{n}+\log\left(1-\frac{p}{n}\right)+o(1)\quad\textnormal{almost surely}.$ (4.2)
Similarly, one can show that
$\log|\hat{\boldsymbol{\Sigma}}^{\textnormal{cen}}_{1:\lfloor nt\rfloor}|$ $=\log\left|\hat{\boldsymbol{\Sigma}}_{1:\lfloor nt\rfloor}\right|+\frac{p}{\lfloor nt\rfloor}+\log\Big(1-\frac{p}{\lfloor nt\rfloor}\Big)+o(1)$ (4.3)
$\log|\hat{\boldsymbol{\Sigma}}^{\textnormal{cen}}_{(\lfloor nt\rfloor+1):n}|$ $=\log\left|\hat{\boldsymbol{\Sigma}}_{(\lfloor nt\rfloor+1):n}\right|+\frac{p}{n-\lfloor nt\rfloor}+\log\Big(1-\frac{p}{n-\lfloor nt\rfloor}\Big)+o(1)$ (4.4)
almost surely. Combining (4.2), (4.3) and (4.4), we can derive a representation of $\log\Lambda_{n,t}^{\textnormal{cen}}$ in terms of $\log\Lambda_{n,t}$, that is
$\frac{2}{n}\log\Lambda_{n,t}^{\textnormal{cen}}$ $=\frac{\lfloor nt\rfloor}{n}\log\left|\hat{\boldsymbol{\Sigma}}^{\textnormal{cen}}_{1:\lfloor nt\rfloor}\right|+\frac{n-\lfloor nt\rfloor}{n}\log\left|\hat{\boldsymbol{\Sigma}}^{\textnormal{cen}}_{(\lfloor nt\rfloor+1):n}\right|-\log\left|\hat{\boldsymbol{\Sigma}}^{\textnormal{cen}}\right|$
$=\frac{\lfloor nt\rfloor}{n}\log\left|\hat{\boldsymbol{\Sigma}}_{1:\lfloor nt\rfloor}\right|+\frac{n-\lfloor nt\rfloor}{n}\log\left|\hat{\boldsymbol{\Sigma}}_{(\lfloor nt\rfloor+1):n}\right|-\log\left|\hat{\boldsymbol{\Sigma}}\right|+\log\Big(1-\frac{p}{\lfloor nt\rfloor}\Big)$
$\quad+\log\Big(1-\frac{p}{n-\lfloor nt\rfloor}\Big)-\log\Big(1-\frac{p}{n}\Big)+\frac{p}{n}+o(1)$
$=\frac{2}{n}\log\Lambda_{n,t}+\log\Big(1-\frac{p}{\lfloor nt\rfloor}\Big)+\frac{n-\lfloor nt\rfloor}{n}\log\Big(1-\frac{p}{n-\lfloor nt\rfloor}\Big)$
$\quad-\frac{\lfloor nt\rfloor}{n}\log\left(1-\frac{p}{n}\right)+\frac{p}{n}+o(1).$ (4.5)
Next, we find a more handy form for the centering term of $\log\Lambda_{n,t}^{\textnormal{cen}}$. As a preparation, we note that
$\left(n-p-\frac{3}{2}\right)\left(\log\left(1-\frac{p}{n}\right)-\log\left(1-\frac{p}{n-1}\right)\right)=\frac{p}{n}+o(1),$
which follows by a Taylor expansion. Then, we calculate
$\frac{\mu_{n,t}}{n}+\log\Big(1-\frac{p}{\lfloor nt\rfloor}\Big)+\log\Big(1-\frac{p}{n-\lfloor nt\rfloor}\Big)-\log\left(1-\frac{p}{n}\right)+\frac{p}{n}$
$=\left(n-p-\frac{3}{2}\right)\log\left(1-\frac{p}{n}\right)-\frac{\lfloor nt\rfloor}{n}\left(\lfloor nt\rfloor-p-\frac{3}{2}\right)\log\left(1-\frac{p}{\lfloor nt\rfloor}\right)+\frac{(\hat{\kappa}_{n}-4)p}{2}$
$\quad-\frac{n-\lfloor nt\rfloor}{n}\left(n-\lfloor nt\rfloor-p-\frac{3}{2}\right)\log\Big(1-\frac{p}{n-\lfloor nt\rfloor}\Big)+\frac{p}{n}$
$=$ $\left(n-p-\frac{3}{2}\right)\log\Big(1-\frac{p}{n-1}\Big)-\frac{\lfloor nt\rfloor}{n}\left(\lfloor nt\rfloor-p-\frac{3}{2}\right)\log\Big(1-\frac{p}{\lfloor nt\rfloor-1}\Big)$
$\quad-\frac{n-\lfloor nt\rfloor}{n}\left(n-\lfloor nt\rfloor-p-\frac{3}{2}\right)\log\Big(1-\frac{p}{n-\lfloor nt\rfloor-1}\Big)+\frac{(\hat{\kappa}_{n}-4)p}{2}+o(1)$
$=\tilde{\mu}_{n,t}+o(1),$ (4.6)
where we note for later usage that the $o(1)$-term does not depend on $t\in[t_{0},1-t_{0}].$ By Theorem 2, (4.5) and (4.6), if follows that for all fixed $k\in\mathbb{N}$, $t_{1},\ldots,t_{k}\in[t_{0},1-t_{0}]$
$\Big(\frac{2\log\Lambda_{n,t_{i}}^{\textnormal{cen}}-\tilde{\mu}_{n,t_{i}}}{n}\Big)_{1\leq i\leq k}\stackrel{{\scriptstyle\mathcal{D}}}{{\to}}(Z(t_{i}))_{1\leq i\leq k}.$ (4.7)
Next, we aim to show that to show that
$\Big(\frac{2\log\Lambda_{n,t}^{\textnormal{cen}}-\tilde{\mu}_{n,t}}{n}\Big)_{t\in[t_{0},1-t_{0}],n\in\mathbb{N}}$ (4.8)
is asymptotically tight. Note that
$\sup_{t\in[t_{0},1-t_{0}]}\Big|2\frac{\log\Lambda_{n,t}^{\textnormal{cen}}-\tilde{\mu}_{n,t}}{n}-2\frac{\log\Lambda_{n,t}-\mu_{n,t}}{n}\Big|$
$=\sup_{t\in[t_{0},1-t_{0}]}\Big|\log\Big(1-\frac{p}{\lfloor nt\rfloor}\Big)+\log\Big(1-\frac{p}{n-\lfloor nt\rfloor}\Big)-\log\left(1-\frac{p}{n}\right)+\frac{p}{n}\Big|+o(1)\lesssim 1$ (4.9)
almost surely. By Theorem 3 and (4.9), we conclude that (4.8) is asymptotically tight. Combining this with (4.7), it follows from Theorem 1.5.4 on Van Der Vaart and Wellner (1996) that
$\Big(\frac{2\log\Lambda_{n,t}^{\textnormal{cen}}-\tilde{\mu}_{n,t}}{n}\Big)_{t\in[t_{0},1-t_{0}]}\stackrel{{\scriptstyle\mathcal{D}}}{{\to}}\big(Z(t)\big)_{t\in[t_{0},1-t_{0}]}\quad\textnormal{in }\ell^{\infty}([t_{0},1-t_{0}]).$
The proof of Theorem 1 concludes by an application of the continuous mapping theorem. ∎
4.1.1 Proof of Theorem 2 - weak convergence of finite-dimensional distributions
In the following, we prove Theorem 2, and the necessary auxiliary results are stated in Section A.1.
Proof of Theorem 2.
For the sake of convenience, we restrict ourselves to the case $k=2.$ Then, using the Cramér–Wold theorem, it suffices to show that
$a_{1}\frac{2\log\Lambda_{n,t_{1}}-\mu_{n,t_{1}}}{n}+a_{2}\frac{2\log\Lambda_{n,t_{2}}-\mu_{n,t_{2}}}{n}\stackrel{{\scriptstyle\mathcal{D}}}{{\to}}\mathcal{N}(0,\tau_{t_{1},t_{2}}^{2})$
for $a_{1},a_{2}\in\mathbb{R},$ where $\tau_{t_{1},t_{2}}^{2}=\operatorname{Var}(a_{1}Z(t_{1})+a_{2}Z(t_{2})).$ In the following, we establish a useful representation of $2\log\Lambda_{n,t}$ by applying a QR-decomposition to several (sub)data matrices. For this purpose, we define for $1\leq i\leq j\leq n$ the matrices
$\mathbf{X}_{i:j}$ $=(\mathbf{x}_{i},\ldots,\mathbf{x}_{j})=(\mathbf{b}_{1,i:j},\ldots,\mathbf{b}_{p,i:j})^{\top}_{\in\mathbb{R}^{p\times(i-j+1)}},\quad\hat{\mathbf{I}}_{i:j}=\frac{1}{j-i+1}\mathbf{X}_{i:j}\mathbf{X}_{i:j}^{\top},\quad$
$\hat{\mathbf{I}}$ $=\hat{\mathbf{I}}_{1:n},\quad\mathbf{b}_{i}=\mathbf{b}_{i,1:n}.$
Moreover, let $\mathbf{P}(i;j:k)$ for $1\leq i\leq p$ and $1\leq j\leq k\leq n$ denote the projection matrix on the orthogonal complement of
$\operatorname{span}\{\mathbf{b}_{1,j:k},\ldots,\mathbf{b}_{i,j:k}\},$
that is, if we let $\mathbf{X}_{i,j:k}=(\mathbf{b}_{1,j:k},\ldots,\mathbf{b}_{i,j:k})^{\top}_{\in\mathbb{R}^{i\times(k-j+1)}},$ then
$\mathbf{P}(i;j:k)=\mathbf{I}-\mathbf{X}_{i,j:k}^{\top}\left(\mathbf{X}_{i,j:k}\mathbf{X}_{i,j:k}^{\top}\right)^{-1}\mathbf{X}_{i,j:k}$
Note that $X_{i:j}=X_{p:i,j}$, set $\mathbf{P}(0;j:k)=\mathbf{I}$ and $\mathbf{P}(i;1:n)=\mathbf{P}(i).$ Before rewriting $\log\Lambda_{n,t}$, we need some preparations. By applying QR-decompositions to $\mathbf{X}_{1:\lfloor nt\rfloor}^{\top},~\mathbf{X}_{(\lfloor nt\rfloor+1):n}^{\top}$ and $\mathbf{X}_{n}^{\top}$ (see (Wang et al., 2018, Section 2) for more details), respectively, we have
[…]
$\left|\lfloor nt\rfloor\hat{\mathbf{I}}_{1:\lfloor nt\rfloor}\right|$ $=\prod_{i=1}^{p}\mathbf{b}_{i,1:\lfloor nt\rfloor}^{\top}\mathbf{P}(i-1;1:\lfloor nt\rfloor)\mathbf{b}_{i,1:\lfloor nt\rfloor},\quad$ (4.10)
$\left|(n-\lfloor nt\rfloor)\hat{\mathbf{I}}_{(\lfloor nt\rfloor+1):n}\right|$ $=\prod_{i=1}^{p}\mathbf{b}_{i,(\lfloor nt\rfloor+1):n}^{\top}\mathbf{P}(i-1;(\lfloor nt\rfloor+1):n)\mathbf{b}_{i,(\lfloor nt\rfloor+1):n}.$ (4.11)
Thus, under the null hypothesis of no change point, the likelihood ratio statistic does not depend on $\boldsymbol{\Sigma}$ and we may write
$2\log\Lambda_{n,t}$ $=2\log\frac{\left|\hat{\mathbf{I}}_{1:\lfloor nt\rfloor}\right|^{\frac{1}{2}\lfloor nt\rfloor}\left|\hat{\mathbf{I}}_{(\lfloor nt\rfloor+1):n}\right|^{\frac{1}{2}(n-\lfloor nt\rfloor)}}{\left|\hat{\mathbf{I}}\right|^{\frac{1}{2}n}}$
[…]
$\quad+np\log n-\lfloor nt\rfloor p\log\lfloor nt\rfloor-(n-\lfloor nt\rfloor)p\log(n-\lfloor nt\rfloor).$
$=\lfloor nt\rfloor\sum_{i=1}^{p}\log\mathbf{b}_{i,1:\lfloor nt\rfloor}^{\top}\mathbf{P}(i-1;1:\lfloor nt\rfloor)\mathbf{b}_{i,1:\lfloor nt\rfloor}$
$+(n-\lfloor nt\rfloor)\sum_{i=1}^{p}\log\mathbf{b}_{i,(\lfloor nt\rfloor+1):n}^{\top}\mathbf{P}(i-1;(\lfloor nt\rfloor+1):n)\mathbf{b}_{i,(\lfloor nt\rfloor+1):n}$
$-n\sum_{i=1}^{p}\log\mathbf{b}_{i}^{\top}\mathbf{P}(i-1)\mathbf{b}_{i}+np\log n-\lfloor nt\rfloor p\log\lfloor nt\rfloor-(n-\lfloor nt\rfloor)p\log(n-\lfloor nt\rfloor).$ (4.13)
Next, we define for $1\leq i\leq p$ and $t\in\{t_{1},t_{2}\}$
$X_{i}$ $=\frac{\mathbf{b}_{i}^{\top}\mathbf{P}(i-1)\mathbf{b}_{i}-(n-i+1)}{n-i+1},$ (4.14)
$X_{i,1:\lfloor nt\rfloor}$ $=\frac{\mathbf{b}_{i,1:\lfloor nt\rfloor}^{\top}\mathbf{P}(i-1;1:\lfloor nt\rfloor)\mathbf{b}_{i,1:\lfloor nt\rfloor}-(\lfloor nt\rfloor-i+1)}{\lfloor nt\rfloor-i+1},$ (4.15)
$X_{i,(\lfloor nt\rfloor+1):n}$ $=\frac{\mathbf{b}_{i,(\lfloor nt\rfloor+1):n}^{\top}\mathbf{P}(i-1;(\lfloor nt\rfloor+1):n)\mathbf{b}_{i,(\lfloor nt\rfloor+1):n}-(n-\lfloor nt\rfloor-i+1)}{n-\lfloor nt\rfloor-i+1},$
$Y_{i}$ $=\log(1+X_{i})-\Big(X_{i}-\frac{X_{i}^{2}}{2}\Big),$ (4.16)
$Y_{i,j:k}$ $=\log(1+X_{i,j:k})-\Big(X_{i,j:k}-\frac{X_{i,j:k}^{2}}{2}\Big),\quad 1\leq j\leq k\leq n.$ (4.17)
Using Stirling’s formula
$\log n!=n\log n-n+\frac{1}{2}\log(2\pi n)+\frac{1}{12n}+\mathcal{O}\left(n^{-3}\right),\quad n\to\infty,$ (4.18)
a straightforward calculation gives
$\sum_{i=1}^{p}\lfloor nt\rfloor\log(\lfloor nt\rfloor-i+1)+\sum_{i=1}^{p}(n-\lfloor nt\rfloor)\log(n-\lfloor nt\rfloor-i+1)-n\sum_{i=1}^{p}\log(n-i+1)$
$\quad+np\log n-\lfloor nt\rfloor p\log\lfloor nt\rfloor-(n-\lfloor nt\rfloor)p\log(n-\lfloor nt\rfloor)$
$=\mu_{n,t}+\frac{n\breve{\sigma}_{n,t}^{2}}{2}+o(n),\quad n\to\infty,$ (4.19)
where
$n\breve{\sigma}_{n,t}^{2}=2n\log\left(1-\frac{p}{n}\right)-2\lfloor nt\rfloor\log\Big(1-\frac{p}{\lfloor nt\rfloor}\Big)-2(n-\lfloor nt\rfloor)\log\Big(1-\frac{p}{n-\lfloor nt\rfloor}\Big)-(\hat{\kappa}_{n}-3)p.$ (4.20)
Combining (4.13) and (4.19) gives the representation
$a_{1}\left(2\log\Lambda_{n,t_{1}}-\mu_{n,t_{1}}\right)+a_{2}\left(2\log\Lambda_{n,t_{2}}-\mu_{n,t_{2}}\right)$
$=\sum_{j=1,2}\Bigg\{a_{j}\sum_{i=1}^{p}\lfloor nt_{j}\rfloor X_{i,1:\lfloor nt_{j}\rfloor}+a_{j}\sum_{i=1}^{p}(n-\lfloor nt_{j}\rfloor)X_{i,(\lfloor nt_{j}\rfloor+1):n}-a_{j}n\sum_{i=1}^{p}X_{i}$
$\quad-a_{j}\left(\sum_{i=1}^{p}\lfloor nt_{j}\rfloor\frac{X_{i,1:\lfloor nt_{j}\rfloor}}{2}^{2}+(n-\lfloor nt_{j}\rfloor)\frac{X_{i,(\lfloor nt_{j}\rfloor+1):n}^{2}}{2}-n\sum_{i=1}^{p}\frac{X_{i}^{2}}{2}-\frac{n\breve{\sigma}_{n,t_{j}}^{2}}{2}\right)$
$\quad+a_{j}\sum_{i=1}^{p}\lfloor nt_{j}\rfloor Y_{i,1:\lfloor nt_{j}\rfloor}+a_{j}\sum_{i=1}^{p}(n-\lfloor nt_{j}\rfloor)Y_{i,(\lfloor nt_{j}\rfloor+1):n}-a_{j}n\sum_{i=1}^{p}Y_{i}\Bigg\}+o(n)$
$=\sum_{j=1,2}\Bigg\{a_{j}\sum_{i=1}^{p}\lfloor nt_{j}\rfloor X_{i,1:\lfloor nt_{j}\rfloor}+a_{j}\sum_{i=1}^{p}(n-\lfloor nt_{j}\rfloor)X_{i,(\lfloor nt_{j}\rfloor+1):n}-a_{j}n\sum_{i=1}^{p}X_{i}\Bigg\}+o_{\mathbb{P}}(n),$
where we applied Lemma 2 and Lemma 1 for the last estimate, which are given in Section A.1. Defining
$D_{i}$ $=\sum_{j=1,2}a_{j}D_{i,j},$
$D_{i,j}$ $=\lfloor nt_{j}\rfloor X_{i,1:\lfloor nt_{j}\rfloor}+(n-\lfloor nt_{j}\rfloor)X_{i,(\lfloor nt_{j}\rfloor+1):n}-nX_{i},\quad 1\leq i\leq p,$ (4.21)
it remains to show that
$\frac{1}{n}\sum_{i=1}^{p}D_{i}\stackrel{{\scriptstyle\mathcal{D}}}{{\to}}\mathcal{N}(0,\tau_{t_{1},t_{2}}^{2}).$ (4.22)
Note that $(D_{i}/n)_{1\leq i\leq p}$ forms a martingale difference scheme with respect to filtration $(\mathcal{A}_{i})_{1\leq i\leq p}$, where the $\sigma$-field $\mathcal{A}_{i}$ is generated by the random variables $\mathbf{b}_{1},\ldots,\mathbf{b}_{i}$ for $1\leq i\leq p.$ In the following, we will show that
$\sum_{i=1}^{p}\mathbb{E}\left[\frac{D_{i,1}D_{i,2}}{n^{2}}\Big|\mathcal{A}_{i-1}\right]$ $=\operatorname{cov}(Z(t_{1}),Z(t_{2}))+o_{\mathbb{P}}(1),$ (4.23)
$\sum_{i=1}^{p}\mathbb{E}\left[D_{i,j}^{2}I\{|D_{i,,j}|>\varepsilon\}\right]$ $=o_{\mathbb{P}}(1),\quad j=1,2,$ (4.24)
By the CLT for martingale differences (see, for example, Corollary 3.1 in Hall and Heyde, 1980), these statements imply (4.22). Regarding (4.24), we have, by Lemma B.26 in Bai and Silverstein (2010), for $\varepsilon>0$
$\mathbb{E}\left|\sum_{i=1}^{p}\mathbb{E}[X_{i,1:\lfloor nt\rfloor}^{2}I\{|X_{i,1:\lfloor nt\rfloor}|>\varepsilon\}|\mathcal{A}_{i-1}]\right|$ $\lesssim\frac{1}{\varepsilon^{\delta/2}}\sum_{i=1}^{p}\mathbb{E}\left|X_{i,1:\lfloor nt\rfloor}\right|^{2+\delta/2}$
$\lesssim\sum_{i=1}^{p}\frac{1}{(\lfloor nt\rfloor-i+1)^{1+\delta/4}}=o(1).$
The other terms in $D_{i,j}$ can be bounded similarly and we (4.24) follows. Next we concentrate on the calculation of the covariance kernel in (4.23). We define for $1\leq j_{1}\leq j_{2}\leq k_{2}\leq k_{1}$ (such that $k_{l}-j_{l}-p>0$ for $l=1,2$
$\mathbf{P}^{j_{2}:k_{2}}(i-1;j_{1}:k_{1})=\left(\left(\mathbf{P}(i-1;j_{1}:k_{1})\right)_{k,l}\right)_{j_{2}\leq k,l\leq k_{2}}\in\mathbb{R}^{(k_{2}-j_{2}+1)\times(k_{2}-j_{2}+1)}.$ (4.25)
In particular, we have $\mathbf{P}^{j_{1}:k_{1}}(i-1;j_{1}:k_{1})=\mathbf{P}(i-1;j_{1}:k_{1})$ and
$\operatorname{tr}\left(\mathbf{P}^{j_{2}:k_{2}}(i-1;j_{1}:k_{1})\mathbf{P}(i-1;j_{2}:k_{2})\right)=(k_{2}-j_{2}-i+1).$
Using formula (9.8.6) in Bai and Silverstein (2010) we calculate for integers $j_{1},j_{2},k_{1},k_{2}$ such that $(k_{1}\wedge k_{2})-(j_{1}\vee j_{2})-p>0$ for $l=1,2$
$n^{2}\sigma^{2}(j_{1},k_{1},j_{2},k_{2}):=\sum_{i=1}^{p}(k_{1}-j_{1}+1)(k_{2}-j_{2}+1)\mathbb{E}\left[X_{i,j_{1}:k_{1}}X_{i,j_{2}:k_{2}}\big|\mathcal{A}_{i-1}\right]$
$=\sum_{i=1}^{p}\frac{(k_{1}-j_{1}+1)(k_{2}-j_{2}+1)}{(k_{1}-j_{1}-i+1)(k_{2}-j_{2}-i+1)}$ (4.26)
$\quad\quad\times\mathbb{E}\Big[\prod_{l=1,2}\left\{\mathbf{b}_{i,j_{l}:k_{l}}^{\top}\mathbf{P}(i-1;j_{l}:k_{l})\mathbf{b}_{i,j_{l}:k_{l}}-(k_{l}-j_{l}-i+1)\right\}\big|\mathcal{A}_{i-1}\Big]$
$=n^{2}\sigma_{1}^{2}(j_{1},k_{1},j_{2},k_{2})+(\mathbb{E}[x_{11}^{4}]-3)n^{2}\sigma_{2}^{2}(j_{1},k_{1},j_{2},k_{2}),$ (4.27)
where
$n^{2}\sigma_{1}^{2}(j_{1},k_{1},j_{2},k_{2})$ $=2\sum_{i=1}^{p}\frac{(k_{1}-j_{1}+1)(k_{2}-j_{2}+1)}{(k_{1}-j_{1}-i+1)\vee(k_{2}-j_{2}-i+1)}$
$\quad\quad\times\operatorname{tr}\left(\mathbf{P}^{(j_{1}\vee j_{2}):(k_{1}\wedge k_{2})}(i-1;j_{1}:k_{1})\mathbf{P}^{(j_{1}\vee j_{2}):(k_{1}\wedge k_{2})}(i-1;j_{2}:k_{2})\right),$
$n^{2}\sigma_{2}^{2}(j_{1},k_{1},j_{2},k_{2})$ $=\sum_{i=1}^{p}\frac{(k_{1}-j_{1}+1)(k_{2}-j_{2}+1)}{(k_{1}-j_{1}-i+1)(k_{2}-j_{2}-i+1)}$
$\quad\quad\times\operatorname{tr}\left(\mathbf{P}^{(j_{1}\vee j_{2}):(k_{1}\wedge k_{2})}(i-1;j_{1}:k_{1})\odot\mathbf{P}^{(j_{1}\vee j_{2}):(k_{1}\wedge k_{2})}(i-1;j_{2}:k_{2})\right)$
and $^{\prime}\odot^{\prime}$ denotes the Hadamard product. We will evaluate these expressions in the case, where $k_{1}$ and $k_{2}$ (and maybe also $j_{1},j_{2}$) are proportional to $n$ using the expansion for the partial sums of the harmonic series
$\sum_{k=1}^{n}\frac{1}{k}=\log n+\gamma+\mathcal{O}\left(\frac{1}{n}\right),\quad n\to\infty,$
(where $\gamma$ denotes the Euler-Mascheroni constant). Using this estimate and (4.25), we obtain for $k_{2}-j_{2}+1=(k_{1}-j_{1}+1)\vee(k_{2}-j_{2}+1)$
$\sigma_{1}^{2}(j_{1},k_{1},j_{2},k_{2})$ $=2\sum_{i=1}^{p}\frac{(k_{1}-j_{1}+1)(k_{2}-j_{2}+1)}{n^{2}(k_{2}-j_{2}-i+1)}$
$=2\frac{(k_{1}-j_{1}+1)(k_{2}-j_{2}+1)}{n^{2}}\sum_{i=k_{2}-j_{2}-p+1}^{k_{2}-j_{2}}\frac{1}{i}$
$=2\frac{(k_{1}-j_{1}+1)(k_{2}-j_{2}+1)}{n^{2}}\Big\{\sum_{i=1}^{k_{2}-j_{2}}\frac{1}{i}-\sum_{i=1}^{k_{2}-j_{2}-p}\frac{1}{i}\Big\}$
$=-2\frac{(k_{1}-j_{1}+1)(k_{2}-j_{2}+1)}{n^{2}}\log\Big(1-\frac{p}{k_{2}-j_{2}}\Big)+o(1)$
$=-2\frac{(k_{1}-j_{1}+1)(k_{2}-j_{2}+1)}{n^{2}}\log\Big(1-\frac{p}{(k_{1}-j_{1})\vee(k_{2}-j_{2})}\Big)+o(1).$ (4.28)
For later use, we note that the $o(1)$ term in (4.28) does not depend on $t\in[t_{0},1-t_{0}]$, if we set $j_{1}=j_{2}=1,k_{1}=k_{2}=\lfloor nt\rfloor$ or $j_{1}=j_{2}=\lfloor nt\rfloor+1,k_{1}=k_{2}=n.$ Moreover, in the case $k_{2}-j_{2}+1=(k_{1}-j_{1}+1)\vee(k_{2}-j_{2}+1)$, it follows from Lemma 3 in Section A.1 below that
$\sigma_{2}^{2}(j_{1},k_{1},j_{2},k_{2})$ $=y\frac{k_{1}-j_{1}+1}{n}+o_{\mathbb{P}}(1),$
$\sigma_{2}^{2}(\lfloor nt_{1}\rfloor+1,n,1,\lfloor nt_{2}\rfloor)$ $=y(t_{2}-t_{1})+o_{\mathbb{P}}(1).$ (4.29)
To calculate $\operatorname{cov}(Z(t_{1}),Z(t_{2}))$ using (4.28), we use that $\sigma^{2}(j_{1},k_{1},j_{2},k_{2})=0$ if $1\leq j_{1}\leq k_{1}<j_{2}\leq k_{2}\leq n$ (this corresponds to the case that $X_{i,j_{1}:k_{1}}$ and $X_{i,j_{2}:k_{2}}$ are independent and thus, for all $1\leq i\leq p$, $\mathbb{E}[X_{i,j_{1}:k_{1}}X_{i,j_{2}:k_{2}}|\mathcal{A}_{i-1}]=0$). In the following, we assume that $t_{1}<t_{2}$, which implies $\sigma^{2}(1,\lfloor nt_{1}\rfloor,\lfloor nt_{2}\rfloor+1,n)=0$. Combining (4.27) and (4.29) gives
$\sum_{i=1}^{p}\mathbb{E}\left[\frac{D_{i,1}D_{i,2}}{n^{2}}\Big|\mathcal{A}_{i-1}\right]$
$=-\sigma^{2}(1,\lfloor nt_{1}\rfloor,1,n)+\sigma^{2}(1,\lfloor nt_{1}\rfloor,1,\lfloor nt_{2}\rfloor)+\sigma^{2}(1,\lfloor nt_{1}\rfloor,\lfloor nt_{2}\rfloor+1,n)$
$\quad-\sigma^{2}(\lfloor nt_{1}\rfloor+1,n,1,n)+\sigma^{2}(\lfloor nt_{1}\rfloor+1,n,1,\lfloor nt_{2}\rfloor)+\sigma^{2}(\lfloor nt_{1}\rfloor+1,n,\lfloor nt_{2}\rfloor+1,n)$
$\quad+\sigma^{2}(1,n,1,n)-\sigma^{2}(1,\lfloor nt_{2}\rfloor,1,n)-\sigma^{2}(\lfloor nt_{2}\rfloor+1,n,1,n)+o_{\mathbb{P}}(1)$
$=-\sigma_{1}^{2}(1,\lfloor nt_{1}\rfloor,1,n)+\sigma_{1}^{2}(1,\lfloor nt_{1}\rfloor,1,\lfloor nt_{2}\rfloor)-\sigma_{1}^{2}(\lfloor nt_{1}\rfloor+1,n,1,n)$
$\quad+\sigma_{1}^{2}(\lfloor nt_{1}\rfloor+1,n,1,\lfloor nt_{2}\rfloor)+\sigma_{1}^{2}(\lfloor nt_{1}\rfloor+1,n,\lfloor nt_{2}\rfloor+1,n)$
$\quad+\sigma_{1}^{2}(1,n,1,n)-\sigma_{1}^{2}(1,\lfloor nt_{2}\rfloor,1,n)-\sigma_{1}^{2}(\lfloor nt_{2}\rfloor+1,n,1,n)+o_{\mathbb{P}}(1).$
Here, we used (4.29) to see that the contributions of the $\sigma_{2}^{2}$-terms cancel each other out. Next, we use Lemma 4 in Section A.1 below to compute the term $\sigma_{1}^{2}(\lfloor nt_{1}\rfloor+1,n,1,\lfloor nt_{2}\rfloor)$. For all remaining $\sigma_{1}^{2}$-terms, we use (4.28) and obtain
$\sum_{i=1}^{p}\mathbb{E}\left[\frac{D_{i,1}D_{i,2}}{n^{2}}\Big|\mathcal{A}_{i-1}\right]$
$=2t_{1}\log(1-y)-2t_{1}t_{2}\log\Big(1-y/t_{2}\Big)+2(1-t_{1})\log(1-y)$
$\quad-2(1-t_{1})t_{2}\log\left(1-\frac{(t_{2}-t_{1})y}{(1-t_{1})t_{2}}\right)-2(1-t_{1})(1-t_{2})\log(1-y/(1-t_{1}))$
$\quad-2\log(1-y)+2t_{2}\log(1-y)+2(1-t_{2})\log(1-y)$
$=2\log(1-y)-2t_{1}t_{2}\log(1-y/t_{2})-2(1-t_{1})t_{2}\log\Big(1-\frac{(t_{2}-t_{1})y}{(1-t_{1})t_{2}}\Big)$
$\quad-2(1-t_{1})(1-t_{2})\log(1-y/(1-t_{1}))+o_{\mathbb{P}}(1)$
$=\operatorname{cov}(Z(t_{1}),Z(t_{2}))+o_{\mathbb{P}}(1).$
If $t_{1}=t_{2}=t$ , then we get
$\operatorname{Var}(Z(t))=2\log(1-y)-2t^{2}\log(1-y/t)-2(1-t)^{2}\log(1-y/(1-t)).$
{funding}
This work was partially supported by the DFG Research unit 5381 Mathematical Statistics in the Information Age, project number 460867398, and the Aarhus University Research Foundation (AUFF), project numbers 47221 and 47388.
[…]
The online supplement contains the proofs of Theorem 1 and its auxiliary results.
A Supplementary Material
A.1 Auxiliary results for the proof of Theorem 2
The convergence of the finite-dimensional distributions is facilitated by the following auxiliary results, whose proofs are postponed to Section A.3. To begin with, we have a result on the quadratic term appearing in the expansion of the test statistic.
Lemma 1.
As $n\to\infty,$ it holds for $t\in[t_{0},1-t_{0}]$
$\sum_{i=1}^{p}\frac{\lfloor nt\rfloor}{n}\frac{X_{i,1:\lfloor nt\rfloor}}{2}^{2}+\frac{n-\lfloor nt\rfloor}{n}\sum_{i=1}^{p}\frac{X_{i,(\lfloor nt\rfloor+1):n}^{2}}{2}-\sum_{i=1}^{p}\frac{X_{i}^{2}}{2}-\frac{\breve{\sigma}_{n,t}^{2}}{2}=o_{\mathbb{P}}(1),$
where $\breve{\sigma}^{2}_{n,t}$ is defined in (4.20).
The following result shows that the logarithmic terms are negligible at a $\delta$-dependent rate. It will also be used in Section A.2 when the proving the asymptotic tightness given in Theorem 3.
Lemma 2.
Assume that (A-1) and (A-2) with some $\delta>0$ are satisfied. Then, it holds for all $t\in[t_{0},1-t_{0}]$
$\frac{1}{n}\sum_{i=1}^{p}\left(\mathbb{E}\left|\lfloor nt\rfloor Y_{i,1:\lfloor nt\rfloor}\right|+\mathbb{E}\left|(n-\lfloor nt\rfloor)Y_{i,(\lfloor nt\rfloor+1):n}\right|+\mathbb{E}\left|nY_{i}\right|\right)\lesssim\frac{1}{n^{\delta/4}},$
where the upper bound does not depend on $t$ and the random variables $Y_{i}$ and $Y_{i,(\lfloor nt\rfloor+1):n}$ are defined in (4.16) and (4.17), respectively.
In the following lemma, we provide an approximation for $\sigma_{2}^{2}$ appearing in (4.29).
Lemma 3.
Suppose that $p<j_{1}\leq j_{2}\leq k_{1}\leq k_{2}\leq n$ such that $k_{2}-j_{2}+1=(k_{1}-j_{1}+1)\vee(k_{2}-j_{2}+1)$. It holds
$\sigma_{2}^{2}(j_{1},k_{1},j_{2},k_{2})$ $=y\frac{k_{1}-j_{1}+1}{n}+o_{\mathbb{P}}(1),$
Moreover, we have for $t_{0}\leq t_{1}<t_{2}\leq t_{0}$
$\sigma_{2}^{2}(\lfloor nt_{1}\rfloor+1,n,1,\lfloor nt_{2}\rfloor)$ $=y(t_{2}-t_{1})+o_{\mathbb{P}}(1).$
We conclude this section by an approximation of $\sigma_{1}^{2}$ defined below (4.27).
Lemma 4.
If $t_{1}<t_{2}$, then we have
$\sigma_{1}^{2}(\lfloor nt_{1}\rfloor+1,n,1,\lfloor nt_{2}\rfloor)=-2(1-t_{1})t_{2}\log\left(1-\frac{(t_{2}-t_{1})y}{(1-t_{1})t_{2}}\right)+o_{\mathbb{P}}(1).$
A.2 Proof of Theorem 3 - asymptotic tightness
We need the following auxiliary results, whose proofs are provided in Section A.5. To begin with, we investigate the increments of the contributing random part of $\log\Lambda_{n,t}$, which is shown to satisfy a finite-dimensional CLT in the proof of Theorem 2.
Lemma 5.
Let Assumption (A-1) and (A-2) with some $\delta>0$ be satisfied and let $t_{1},t_{2}\in[t_{0},1-t_{0}]$ and $D_{i,j}$ be defined as in (4.21) for $j\in\{1,2\},~1\leq i\leq p.$ Then, there exists random variables $Z_{1}=Z_{1,n}(t_{1},t_{2}),Z_{2}=Z_{2,n}(t_{1},t_{2})$ such that
$\frac{1}{n}\sum\limits_{i=1}^{p}\left(D_{i,1}-D_{i,2}\right)=Z_{1}+Z_{2}$
and
$\mathbb{E}[Z_{1}^{2}]$ $\lesssim\left|\frac{\lfloor nt_{1}\rfloor-\lfloor nt_{2}\rfloor}{n}\right|^{1+d}$
$\mathbb{E}[|Z_{2}|^{2+\delta/2}]$ $\lesssim\left|\frac{\lfloor nt_{1}\rfloor-\lfloor nt_{2}\rfloor}{n}\right|^{1+d},$
for some $d>0.$
Next, we need a uniform result on the quadratic terms, which is provided in the next lemma.
Lemma 6.
If Assumption (A-1) and (A-2) with some $\delta>4$ are satisfied, then there exist random variables $Q_{n,1,t}$ and $Q_{n,2,t}$ with
$Q_{n,1,t}+Q_{n,2,t}=\sum_{i=1}^{p}\frac{\lfloor nt\rfloor}{n}\frac{X_{i,1:\lfloor nt\rfloor}}{2}^{2}+\frac{n-\lfloor nt\rfloor}{n}\sum_{i=1}^{p}\frac{X_{i,(\lfloor nt\rfloor+1):n}^{2}}{2}-\sum_{i=1}^{p}\frac{X_{i}^{2}}{2}-\frac{\breve{\sigma}_{n,t}^{2}}{2},$ (A.1)
such that $(Q_{n,1,t})$ is asymptotically tight in $\ell^{\infty}([t_{0},1-t_{0}])$ and $(Q_{2,n,t})$ satisfies the moment inequality
$\sup_{t\in[t_{0},1-t_{0}]}\mathbb{E}|Q_{2,n,t}|^{2+\delta/4}\lesssim\frac{1}{n^{1+\delta/8}}.$ (A.2)
Finally, we recall Lemma 2 given in Section A.1 on the logarithmic terms. Using these auxiliary results, we are in the position to give a proof of Theorem 3.
Proof of Theorem 3.
By Lemma 6 and (4.19), it suffices to show that $\{L_{n,t_{1}}\}_{t_{1}\in[t_{0},1-t_{0}]}$ with
$L_{n,t_{1}}:=\frac{1}{n}\sum_{i=1}^{p}D_{i,1}-Q_{2,n,t_{1}}+\sum_{i=1}^{p}\frac{\lfloor nt_{1}\rfloor}{n}Y_{i,1:\lfloor nt_{1}\rfloor}+a_{j}\sum_{i=1}^{p}\frac{n-\lfloor nt_{1}\rfloor}{n}Y_{i,(\lfloor nt_{1}\rfloor+1):n}-\sum_{i=1}^{p}Y_{i}$
is asymptotically tight. We write for $t_{1},t_{2}\in[t_{0},1-t_{0}]$
$L_{n,t_{1}}-L_{n,t_{2}}=Z_{1,n}(t_{1},t_{2})+Z_{2,n}(t_{1},t_{2})+R_{n}(t_{1})+R_{n}(t_{2})-Q_{2,n,t_{1}}+Q_{2,n,t_{2}},$
where $Z_{1,n}(t_{1},t_{2}),Z_{2,n}(t_{1},t_{2})$ are the random variables in Lemma 5, and
$nR_{n}(t_{1})$ $=\sum_{i=1}^{p}\lfloor nt_{1}\rfloor Y_{i,1:\lfloor nt_{1}\rfloor}+\sum_{i=1}^{p}(n-\lfloor nt_{1}\rfloor)Y_{i,(\lfloor nt_{1}\rfloor+1):n}-n\sum_{i=1}^{p}Y_{i}.$
For analyzing the increments of $(L_{n,t})$ we define for $t_{0}\leq r\leq s\leq t\leq 1-t_{0}$
$m(r,s,t)=\min\{|L_{n,s}-L_{n,t}|,|L_{n,r}-L_{n,s}|\}.$
Note that under the moment assumption (A-2) with $\delta>4$, we have by Lemma 2 and Lemma 6
$\sup_{t\in[t_{0},1-t_{0}]}\left(\mathbb{E}|Q_{2,n,t}|^{2+\delta/4}\vee\mathbb{E}|R_{n}(t)|\right)\lesssim\frac{1}{n^{1+d}}$ (A.3)
for some $d>0$, which may be chosen such that it coincides with the $d>0$ from Lemma 5. Note that if tr<1/nt-r<1/n, we have $\lfloor nr\rfloor=\lfloor ns\rfloor$ or $\lfloor ns\rfloor=\lfloor nt\rfloor$, and thus, $m(r,s,t)=0$ almost surely. If $t-r\geq 1/n,$ it holds for all $\lambda>0$ by Lemma 5 and (A.3),
$\mathbb{P}(m(r,s,t)>\lambda)$
$\lesssim\mathbb{E}|Z_{1,n}(s,t)|^{2}+\mathbb{E}|Z_{1,n}(r,s)|^{2}+\mathbb{E}|Z_{2,n}(s,t)|^{2+\delta/2}+\mathbb{E}|Z_{2,n}(r,s)|^{2+\delta/2}$
$\quad+\sup_{t\in[t_{0},1-t_{0}]}\left(\mathbb{E}|Q_{2,n,t}|^{2+\delta/4}+\mathbb{E}|R_{n}(t)|\right)$
$\lesssim\left(\frac{\lfloor nt\rfloor-\lfloor ns\rfloor}{n}\right)^{1+d}+\left(\frac{\lfloor ns\rfloor-\lfloor nr\rfloor}{n}\right)^{1+d}+\frac{1}{n^{1+d}}\lesssim\left(t-r+\frac{1}{n}\right)^{1+d}+(t-r)^{1+d}$
$\lesssim(t-r)^{1+d}.$ (A.4)
Similarly, we get
$\mathbb{P}(|L_{n,t}-L_{n,s}|>\lambda)\lesssim\left(t-s+\frac{1}{n}\right)^{1+d}+\frac{1}{n^{1+d}}.$ (A.5)
Define
$K_{j}=\left[\frac{j-1}{m},\frac{j}{m}\right],\quad\lfloor mt_{0}\rfloor\leq j\leq\lfloor m(1-t_{0})\rfloor,\quad m\in\mathbb{N}.$
Combining (A.4) and (A.5) with Corollary A.4 in Dette and Tomecki (2019), we have for $\lfloor mt_{0}\rfloor\leq j\leq\lfloor m(1-t_{0})\rfloor$
$\mathbb{P}\left(\sup_{t_{1},t_{2}\in K_{j}}\left|L_{n,t_{1}}-L_{n,t_{2}}\right|>\lambda\right)\lesssim\frac{1}{m^{1+d}}+\left(\frac{1}{n}+\frac{1}{m}\right)^{1+d}+\frac{1}{n^{1+d}}.$ (A.6)
This implies
$\limsup_{n\to\infty}\mathbb{P}\left(\sup_{\lfloor mt_{0}\rfloor\leq j\leq m}\sup_{s,t\in K_{j}}|L_{n,t_{1}}-L_{n,t_{2}}|>\lambda\right)\lesssim\frac{1}{m^{d}}\to 0,\textnormal{ as }m\to\infty.$
Since the finite-dimensional distributions of $\boldsymbol{\Lambda}_{n}$ and so, those of $(L_{n,t})$, converge weakly, we conclude from (A.6) and Theorem 1.5.6 in Van Der Vaart and Wellner (1996) that $(L_{n,t})$ is asymptotically tight.
A.3 Auxiliary results
In this section, we provide the proofs of the auxiliary results given in section A.1, among others. Note that the proof of Lemma 2 is very similar to the proof of Lemma 3 in Dörnemann (2023) and we skip it for the sake of brevity. To begin with, we prove Lemma 4 providing an approximation for the quantity $\sigma_{1}^{2}$ defined below (4.27).
Proof of Lemma 4.
Recalling the representation of $\sigma_{1}^{2}$ below (4.27) we obtain
$n^{2}\sigma_{1}^{2}(\lfloor nt_{1}\rfloor+1,n,1,\lfloor nt_{2}\rfloor)):=\sum_{i=1}^{p}(n-\lfloor nt_{1}\rfloor)\lfloor nt_{2}\rfloor\mathbb{E}\left[X_{i,(\lfloor nt_{1}\rfloor+1):n}X_{i,1:\lfloor nt_{2}\rfloor}\big|\mathcal{A}_{i-1}\right]$
$=2\sum_{i=1}^{p}\frac{(n-\lfloor nt_{1}\rfloor)\lfloor nt_{2}\rfloor}{(n-\lfloor nt_{1}\rfloor-i+1)(\lfloor nt_{2}\rfloor-i+1)}$
$\quad\quad\quad\times\operatorname{tr}\left(\mathbf{P}^{(\lfloor nt_{1}\rfloor+1):\lfloor nt_{2}\rfloor}(i-1;(\lfloor nt_{1}\rfloor+1):n)\mathbf{P}^{(\lfloor nt_{1}\rfloor+1):\lfloor nt_{2}\rfloor}(i-1;1:\lfloor nt_{2}\rfloor)\right)$ (A.7)
where
$\operatorname{tr}\left(\mathbf{P}^{(\lfloor nt_{1}\rfloor+1):\lfloor nt_{2}\rfloor}(i-1;(\lfloor nt_{1}\rfloor+1):n)\mathbf{P}^{(\lfloor nt_{1}\rfloor+1):\lfloor nt_{2}\rfloor}(i-1;1:\lfloor nt_{2}\rfloor)\right)$
$=\sum_{k,l=\lfloor nt_{1}\rfloor+1}^{\lfloor nt_{2}\rfloor}\left(\mathbf{P}(i-1;(\lfloor nt_{1}\rfloor+1):n)\right)_{kl}\left(\mathbf{P}(i-1;1:\lfloor nt_{2}\rfloor)\right)_{kl}$
Let
$\mathbf{S}_{i,j:k}=\frac{1}{n}\mathbf{X}_{i,j:k}\mathbf{X}_{i,j:k}^{\top}$
for $1\leq i\leq p,~1\leq j<k\leq p.$ Then, we may write (replacing for a moment $i$ by i1i-1)
$\sum_{k,l=\lfloor nt_{1}\rfloor+1}^{\lfloor nt_{2}\rfloor}\left(\mathbf{P}(i;(\lfloor nt_{1}\rfloor+1):n)\right)_{kl}\left(\mathbf{P}(i;1:\lfloor nt_{2}\rfloor)\right)_{kl}=\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor-S_{i,1}-S_{i,2}+S_{i,3},$ (A.8)
where
$S_{i,1}$ $=\operatorname{tr}\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):\lfloor nt_{2}\rfloor}\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{-1},$
$S_{i,2}$ $=\operatorname{tr}\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):\lfloor nt_{2}\rfloor}\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{-1},$
$S_{i,3}$ $=\operatorname{tr}\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):\lfloor nt_{2}\rfloor}\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{-1}\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):\lfloor nt_{2}\rfloor}\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{-1}.$
In the following, we will approximate the quantities $S_{i,1},S_{i,2}$ and $S_{i,3}.$ Note that these terms actually depend on $n,~t_{1},~t_{2}$, which is not reflected by our notation. Moreover, it is important to emphasize that, for instance, the product $\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):\lfloor nt_{2}\rfloor}\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{-1}$ is not an F-matrix in the classical sense, since the data matrices $\mathbf{X}_{i,(\lfloor nt_{1}\rfloor+1):\lfloor nt_{2}\rfloor}$ and $\mathbf{X}_{i,1:\lfloor nt_{2}\rfloor}$ are dependent.
Calculation of $S_{i,1}$ By an application of the Sherman-Morrison formula, we obtain
$\mathbf{S}_{i,j:k}^{-1}=\left(\mathbf{S}_{i,j:k}^{(-l)}\right)^{-1}-\frac{1}{n}\beta_{i,j:k}^{(-l)}\left(\mathbf{S}_{i,j:k}^{(-l)}\right)^{-1}\mathbf{x}_{i,k}\mathbf{x}_{i,k}^{\top}\left(\mathbf{S}_{i,j:k}^{(-l)}\right)^{-1},\quad 1\leq j\leq l\leq k\leq n,~j\neq k,$ (A.9)
where
$\mathbf{S}_{i,j:k}^{(-l)}$ $=\frac{1}{n}\sum\limits_{m=j}^{k}\mathbf{x}_{i,m}\mathbf{x}_{i,m}^{\top}-\frac{1}{n}\mathbf{x}_{i,l}\mathbf{x}_{i,l}^{\top},$
$\beta_{i,j:k}^{(-l)}$ $=\frac{1}{1+n^{-1}\mathbf{x}_{i,l}^{\top}\left(\mathbf{S}_{i,j:k}^{(-l)}\right)^{-1}\mathbf{x}_{i,l}},$
$\mathbf{x}_{i,l}$ $=(x_{l1},\ldots,x_{li})^{\top}.$
As a preparation, we first calculate the mean of $\beta_{i,j:k}^{(-l)}$. Using the identity (6.1.11) in Bai and Silverstein (2010), we have
$\mathbf{I}_{i}=\frac{1}{n}\sum_{l=j}^{k}\mathbf{x}_{i,l}\mathbf{x}_{i,l}^{\top}\mathbf{S}_{i,j:k}^{-1}=\sum_{l=j}^{k}\frac{\frac{1}{n}\mathbf{x}_{i,l}\mathbf{x}_{i,l}^{\top}\left(\mathbf{S}_{i,j:k}^{(-l)}\right)^{-1}}{1+\frac{1}{n}\mathbf{x}_{i,l}^{\top}\left(\mathbf{S}_{i,j:k}^{(-l)}\right)^{-1}\mathbf{x}_{i,l}}.$
Applying the trace on both sides and dividing by kj+1k-j+1, yields
$i=\sum_{m=j}^{k}\left(1-\beta_{i,j:k}^{(-l)}\right),$
which implies by the i.i.d. assumption,
$\mathbb{E}[\beta_{i,j:k}^{(-l)}]=1-\frac{i}{k-j+1}=\frac{k-j-i+1}{k-j+1}.$ (A.10)
Moreover, note that $\mathbf{S}_{i,j:k}^{-1}||\leq 1/((1-\sqrt{t_{0}})^{2}-\varepsilon)<\infty$ for some $\varepsilon>0$ and all large $n.$ As a further preparation, we note that $(\lfloor nt_{1}\rfloor/n)*(1/i)\operatorname{tr}\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{-1}$ can be approximated by the first negative moment of the Marčenko–Pastur distribution $F^{i/\lfloor nt_{2}\rfloor}$, that is,
$\frac{\lfloor nt_{2}\rfloor}{in}\operatorname{tr}\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{-1}=\frac{1}{1-i/\lfloor nt_{2}\rfloor}+o_{\mathbb{P}}(1),$ (A.11)
uniformly with respect to $1\leq i\leq p$. Using (A.9), (A.10) and Lemma B.26 in Bai and Silverstein (2010), we get for the first term
$S_{i,1}=$ $\operatorname{tr}\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):\lfloor nt_{2}\rfloor}\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{-1}$
$=\frac{1}{n}\sum_{k=\lfloor nt_{1}\rfloor+1}^{\lfloor nt_{2}\rfloor}\mathbf{x}_{i,k}^{\top}\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{-1}\mathbf{x}_{i,k}$
$=\sum_{k=\lfloor nt_{1}\rfloor+1}^{\lfloor nt_{2}\rfloor}\left\{\frac{1}{n}\mathbf{x}_{i,k}^{\top}\left(\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{(-k)}\right)^{-1}\mathbf{x}_{i,k}-\frac{1}{n^{2}}\beta_{i,1:\lfloor nt_{2}\rfloor}^{(-k)}\left(\mathbf{x}_{i,k}^{\top}\left(\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{(-k)}\right)^{-1}\mathbf{x}_{i,k}\right)^{2}\right\}$
$=\sum_{k=\lfloor nt_{1}\rfloor+1}^{\lfloor nt_{2}\rfloor}\left\{\frac{1}{n}\operatorname{tr}\left(\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{(-k)}\right)^{-1}-\frac{1}{n^{2}}(1-i/\lfloor nt_{2}\rfloor)\left(\operatorname{tr}\left(\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{(-k)}\right)^{-1}\right)^{2}\right\}+o_{\mathbb{P}}(n)$
$=\frac{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor}{n}\left\{\operatorname{tr}\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{-1}-\frac{1-i/\lfloor nt_{2}\rfloor}{n}\left(\operatorname{tr}\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{-1}\right)^{2}\right\}+o_{\mathbb{P}}(n).$
Combining this with (A.11), we get
$\frac{1}{n}S_{i,1}$ $=\frac{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor}{n}\left\{\frac{i}{\lfloor nt_{2}\rfloor}\frac{1}{1-\frac{i}{\lfloor nt_{2}\rfloor}}-\left(1-\frac{i}{\lfloor nt_{2}\rfloor}\right)\left(\frac{i}{\lfloor nt_{2}\rfloor}\frac{1}{1-\frac{i}{\lfloor nt_{2}\rfloor}}\right)^{2}\right\}+o_{\mathbb{P}}(1)$
$=\frac{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor}{n}\frac{i}{\lfloor nt_{2}\rfloor}+o_{\mathbb{P}}(1)$ (A.12)
uniformly with respect to $1\leq i\leq p$.
Calculation of $S_{i,2}$ Similarly to the previous step, we may show that
$\frac{1}{n}S_{i,2}=\frac{(\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor)i}{n(n-\lfloor nt_{1}\rfloor)}+o_{\mathbb{P}}(1)$ (A.13)
uniformly with respect to $1\leq i\leq p$.
Calculation of $S_{i,3}$ We decompose $S_{i,3}$ as
$S_{i,3}$ $=\frac{1}{n^{2}}\sum_{k,l=\lfloor nt_{1}\rfloor+1}^{\lfloor nt_{2}\rfloor}\mathbf{x}_{i,l}^{\top}\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{-1}\mathbf{x}_{i,k}\mathbf{x}_{i,k}^{\top}\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{-1}\mathbf{x}_{i,l}=S_{i,3,1}+S_{i,3,2},$
where
$S_{i,3,1}$ $=\frac{1}{n^{2}}\sum_{k=\lfloor nt_{1}\rfloor+1}^{\lfloor nt_{2}\rfloor}\mathbf{x}_{i,k}^{\top}\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{-1}\mathbf{x}_{i,k}\mathbf{x}_{i,k}^{\top}\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{-1}\mathbf{x}_{i,k},$
$S_{i,3,2}$ $=\frac{1}{n^{2}}\sum_{\begin{subarray}{c}k,l=\lfloor nt_{1}\rfloor+1,\\ k\neq l\end{subarray}}^{\lfloor nt_{2}\rfloor}\mathbf{x}_{i,l}^{\top}\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{-1}\mathbf{x}_{i,k}\mathbf{x}_{i,k}^{\top}\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{-1}\mathbf{x}_{i,l}.$
These terms will be further investigated in the following steps.
Calculation of $\mathbf{S}_{i,3,1}$ Applying similar techniques as in the previous steps, we get
$\frac{1}{n}S_{i,3,1}$ $=\frac{1}{n^{3}}\sum_{k=\lfloor nt_{1}\rfloor+1}^{\lfloor nt_{2}\rfloor}\mathbf{x}_{i,k}^{\top}\left(\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{(-k)}\right)^{-1}\mathbf{x}_{i,k}\mathbf{x}_{i,k}^{\top}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{(-k)}\right)^{-1}\mathbf{x}_{i,k}$
$-\frac{1}{n^{4}}\sum_{k=\lfloor nt_{1}\rfloor+1}^{\lfloor nt_{2}\rfloor}\beta_{i,1:\lfloor nt_{2}\rfloor}^{(-k)}\left(\mathbf{x}_{i,k}^{\top}\left(\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{(-k)}\right)^{-1}\mathbf{x}_{i,k}\right)^{2}\mathbf{x}_{i,k}^{\top}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{(-k)}\right)^{-1}\mathbf{x}_{i,k}$
$-\frac{1}{n^{4}}\sum_{k=\lfloor nt_{1}\rfloor+1}^{\lfloor nt_{2}\rfloor}\beta_{i,(\lfloor nt_{1}\rfloor+1):n}^{(-k)}\mathbf{x}_{i,k}^{\top}\left(\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{(-k)}\right)^{-1}\mathbf{x}_{i,k}\left(\mathbf{x}_{i,k}^{\top}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{(-k)}\right)^{-1}\mathbf{x}_{i,k}\right)^{2}$
$+\frac{1}{n^{5}}\sum_{k=\lfloor nt_{1}\rfloor+1}^{\lfloor nt_{2}\rfloor}\beta_{i,1:\lfloor nt_{2}\rfloor}^{(-k)}\beta_{i,(\lfloor nt_{1}\rfloor+1):n}^{(-k)}\left(\mathbf{x}_{i,k}^{\top}\left(\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{(-k)}\right)^{-1}\mathbf{x}_{i,k}\right)^{2}\left(\mathbf{x}_{i,k}^{\top}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{(-k)}\right)^{-1}\mathbf{x}_{i,k}\right)^{2}$
$=$ $\frac{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor}{n}\Bigg\{\frac{1}{n^{2}}\operatorname{tr}\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{-1}\operatorname{tr}\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{-1}-\frac{\lfloor nt_{2}\rfloor-i}{\lfloor nt_{2}\rfloor}\frac{1}{n^{3}}\left(\operatorname{tr}\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{-1}\right)^{2}\operatorname{tr}\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{-1}$
$-\frac{n-\lfloor nt_{1}\rfloor-i}{n-\lfloor nt_{1}\rfloor}\frac{1}{n^{3}}\operatorname{tr}\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{-1}\left(\operatorname{tr}\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{-1}\right)^{2}$
$+\frac{\lfloor nt_{2}\rfloor-i}{\lfloor nt_{2}\rfloor}\frac{n-\lfloor nt_{1}\rfloor-i}{n-\lfloor nt_{1}\rfloor}\frac{1}{n^{4}}\left(\operatorname{tr}\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{-1}\operatorname{tr}\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{-1}\right)^{2}\Bigg\}+o_{\mathbb{P}}(n)$
In the following, we use a general form of (A.11), namely,
$\frac{1}{n}\operatorname{tr}\mathbf{S}_{i,j:k}^{-1}=\frac{i}{k-j-i+1}+o_{\mathbb{P}}(1),\quad 1\leq j<k\leq n,$ (A.14)
uniformly with respect to $1\leq i\leq p.$ This gives
$\frac{1}{n}S_{i,3,1}$ $=\frac{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor}{n}\Bigg\{\frac{i^{2}}{(\lfloor nt_{2}\rfloor-i)(n-\lfloor nt_{1}\rfloor-i)}$
$-\frac{\lfloor nt_{2}\rfloor-i}{\lfloor nt_{2}\rfloor}\frac{i^{3}}{(\lfloor nt_{2}\rfloor-i)^{2}(n-\lfloor nt_{1}\rfloor-i)}$
$-\frac{n-\lfloor nt_{1}\rfloor-i}{n-\lfloor nt_{1}\rfloor}\frac{i^{3}}{(\lfloor nt_{2}\rfloor-i)(n-\lfloor nt_{1}\rfloor-i)^{2}}$
$+\frac{\lfloor nt_{2}\rfloor-i}{\lfloor nt_{2}\rfloor}\frac{n-\lfloor nt_{1}\rfloor-i}{n-\lfloor nt_{1}\rfloor}\frac{i^{4}}{(\lfloor nt_{2}\rfloor-i)^{2}(n-\lfloor nt_{1}\rfloor-i)^{2}}\Bigg\}+o_{\mathbb{P}}(1)$
$=\frac{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor}{n}\Bigg\{\frac{i^{2}}{(\lfloor nt_{2}\rfloor-i)(n-\lfloor nt_{1}\rfloor-i)}-\frac{i^{3}}{\lfloor nt_{2}\rfloor(\lfloor nt_{2}\rfloor-i)(n-\lfloor nt_{1}\rfloor-i)}$
$-\frac{i^{3}}{(n-\lfloor nt_{1}\rfloor)(\lfloor nt_{2}\rfloor-i)(n-\lfloor nt_{1}\rfloor-i)}$
$+\frac{i^{4}}{\lfloor nt_{2}\rfloor(n-\lfloor nt_{1}\rfloor)(\lfloor nt_{2}\rfloor-i)(n-\lfloor nt_{1}\rfloor-i)}\Bigg\}+o_{\mathbb{P}}(1)$
$=\frac{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor}{n}\frac{i^{2}}{\lfloor nt_{2}\rfloor(n-\lfloor nt_{1}\rfloor)}+o_{\mathbb{P}}(1)$ (A.15)
Calculation of $\mathbf{S}_{i,3,2}$ Again applying similar techniques as in the previous steps, especially (A.9), (A.10) and (A.14), we get
$\frac{1}{n}\operatorname{tr}\mathbf{S}_{i,3,2}$
$=\frac{1}{n^{3}}\sum_{\begin{subarray}{c}k,l=\lfloor nt_{1}\rfloor+1,\\ k\neq l\end{subarray}}^{\lfloor nt_{2}\rfloor}\mathbf{x}_{i,l}^{\top}\left(\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{(-l)}\right)^{-1}\mathbf{x}_{i,k}\mathbf{x}_{i,k}^{\top}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{(-l)}\right)^{-1}\mathbf{x}_{i,l}$
$-\frac{1}{n^{4}}\sum_{\begin{subarray}{c}k,l=\lfloor nt_{1}\rfloor+1,\\ k\neq l\end{subarray}}^{\lfloor nt_{2}\rfloor}\beta_{i,1:\lfloor nt_{2}\rfloor}^{(-l)}\mathbf{x}_{i,l}^{\top}\left(\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{(-l)}\right)^{-1}\mathbf{x}_{i,l}\mathbf{x}_{i,l}^{\top}\left(\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{(-l)}\right)^{-1}\mathbf{x}_{i,k}\mathbf{x}_{i,k}^{\top}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{(-l)}\right)^{-1}\mathbf{x}_{i,l}$
$-\frac{1}{n^{4}}\sum_{\begin{subarray}{c}k,l=\lfloor nt_{1}\rfloor+1,\\ k\neq l\end{subarray}}^{\lfloor nt_{2}\rfloor}\beta_{i,(\lfloor nt_{1}\rfloor+1):n}^{(-l)}\mathbf{x}_{i,l}^{\top}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{(-l)}\right)^{-1}\mathbf{x}_{i,l}\mathbf{x}_{i,l}^{\top}\left(\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{(-l)}\right)^{-1}\mathbf{x}_{i,k}\mathbf{x}_{i,k}^{\top}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{(-l)}\right)^{-1}\mathbf{x}_{i,l}$
$+\frac{1}{n^{5}}\sum_{\begin{subarray}{c}k,l=\lfloor nt_{1}\rfloor+1,\\ k\neq l\end{subarray}}^{\lfloor nt_{2}\rfloor}\Big\{\beta_{i,1:\lfloor nt_{2}\rfloor}^{(-l)}\beta_{i,(\lfloor nt_{1}\rfloor+1):n}^{(-l)}\mathbf{x}_{i,l}^{\top}\left(\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{(-l)}\right)^{-1}\mathbf{x}_{i,l}\mathbf{x}_{i,l}^{\top}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{(-l)}\right)^{-1}\mathbf{x}_{i,l}$
$\quad~\quad~\quad~\quad~\quad~\quad\times\mathbf{x}_{i,l}^{\top}\left(\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{(-l)}\right)^{-1}\mathbf{x}_{i,k}\mathbf{x}_{i,k}^{\top}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{(-l)}\right)^{-1}\mathbf{x}_{i,l}\Big\}$
$=\frac{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor}{n^{3}}\sum_{\begin{subarray}{c}k=\lfloor nt_{1}\rfloor+1\end{subarray}}^{\lfloor nt_{2}\rfloor}\mathbf{x}_{i,k}^{\top}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}\right)^{-1}\left(\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}\right)^{-1}\mathbf{x}_{i,k}$
$-\frac{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor}{n^{4}}\beta_{i,1:\lfloor nt_{2}\rfloor}\operatorname{tr}\left(\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}\right)^{-1}\sum_{\begin{subarray}{c}k=\lfloor nt_{1}\rfloor+1\end{subarray}}^{\lfloor nt_{2}\rfloor}\mathbf{x}_{i,k}^{\top}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}\right)^{-1}\left(\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}\right)^{-1}\mathbf{x}_{i,k}$
$-\frac{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor}{n^{4}}\beta_{i,(\lfloor nt_{1}\rfloor+1):n}\operatorname{tr}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}\right)^{-1}\sum_{\begin{subarray}{c}k=\lfloor nt_{1}\rfloor+1\end{subarray}}^{\lfloor nt_{2}\rfloor}\mathbf{x}_{i,k}^{\top}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}\right)^{-1}\left(\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}\right)^{-1}\mathbf{x}_{i,k}$
$+\frac{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor}{n^{5}}\beta_{i,1:\lfloor nt_{2}\rfloor}\beta_{i,(\lfloor nt_{1}\rfloor+1):n}\operatorname{tr}\left(\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}\right)^{-1}\operatorname{tr}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}\right)^{-1}$
$\quad\times\sum_{\begin{subarray}{c}k=\lfloor nt_{1}\rfloor+1\end{subarray}}^{\lfloor nt_{2}\rfloor}\mathbf{x}_{i,k}^{\top}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}\right)^{-1}\left(\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}\right)^{-1}\mathbf{x}_{i,k}$
$+o_{\mathbb{P}}(1)$
$=\frac{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor}{n}\frac{1}{n^{2}}\sum_{\begin{subarray}{c}k=\lfloor nt_{1}\rfloor+1\end{subarray}}^{\lfloor nt_{2}\rfloor}\mathbf{x}_{i,k}^{\top}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}\right)^{-1}\left(\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}\right)^{-1}\mathbf{x}_{i,k}$
$\quad\times\Big\{1+\frac{i^{2}}{\left(n-\lfloor nt_{1}\rfloor\right)\left(\lfloor nt_{2}\rfloor\right)}-\frac{i}{n-\lfloor nt_{1}\rfloor}-\frac{i}{\lfloor nt_{2}\rfloor}\Big\}+o_{\mathbb{P}}(1)$
$=\frac{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor}{n}\frac{\left(n-\lfloor nt_{1}\rfloor-i\right)\left(\lfloor nt_{2}\rfloor-i\right)}{\lfloor nt_{2}\rfloor\left(n-\lfloor nt_{1}\rfloor\right)}\frac{1}{n^{2}}\sum_{\begin{subarray}{c}k=\lfloor nt_{1}\rfloor+1\end{subarray}}^{\lfloor nt_{2}\rfloor}\mathbf{x}_{i,k}^{\top}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}\right)^{-1}\left(\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}\right)^{-1}\mathbf{x}_{i,k}$
$+o_{\mathbb{P}}(1).$ (A.16)
Thus, we need to compute
$\frac{1}{n^{2}}\sum_{\begin{subarray}{c}k=\lfloor nt_{1}\rfloor+1\end{subarray}}^{\lfloor nt_{2}\rfloor}\mathbf{x}_{i,k}^{\top}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}\right)^{-1}\left(\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}\right)^{-1}\mathbf{x}_{i,k}$
$=\frac{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor}{n}\frac{1}{n}\operatorname{tr}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{-1}\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{-1}\right)\Big\{1-\beta_{i,(\lfloor nt_{1}\rfloor+1):n}\frac{1}{n}\operatorname{tr}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{-1}\right)$
$\quad-\beta_{i,1:\lfloor nt_{2}\rfloor}\frac{1}{n}\operatorname{tr}\left(\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{-1}\right)+\beta_{i,1:\lfloor nt_{2}\rfloor}\beta_{i,(\lfloor nt_{1}\rfloor+1):n}\frac{1}{n^{2}}\operatorname{tr}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{-1}\right)\operatorname{tr}\left(\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{-1}\right)\Big\}$
$=\frac{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor}{n}\frac{\left(n-\lfloor nt_{1}\rfloor-i\right)\left(\lfloor nt_{2}\rfloor-i\right)}{\lfloor nt_{2}\rfloor\left(n-\lfloor nt_{1}\rfloor\right)}\frac{1}{n}\operatorname{tr}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{-1}\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{-1}\right)+o_{\mathbb{P}}(1)$ (A.17)
Combining (A.16), (A.17) and Lemma 7, we get
$\frac{1}{n}\operatorname{tr}\mathbf{S}_{i,3,2}$
$=\left(\frac{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor}{n}\frac{\left(n-\lfloor nt_{1}\rfloor-i\right)\left(\lfloor nt_{2}\rfloor-i\right)}{\lfloor nt_{2}\rfloor\left(n-\lfloor nt_{1}\rfloor\right)}\right)^{2}\frac{1}{n}\operatorname{tr}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{-1}\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{-1}\right)$
$=\left(\frac{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor}{n}\right)^{2}\frac{\left(n-\lfloor nt_{1}\rfloor-i\right)\left(\lfloor nt_{2}\rfloor-i\right)}{\lfloor nt_{2}\rfloor\left(n-\lfloor nt_{1}\rfloor\right)}\frac{in}{\left(n-\lfloor nt_{1}\rfloor\right)\lfloor nt_{2}\rfloor-\left(\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor\right)i}$
$+o_{\mathbb{P}}(1)$ (A.18)
Conclusion Using (A.7) and (A.8), we obtain
$\sigma^{2}(\lfloor nt_{1}\rfloor+1,n,1,\lfloor nt_{2}\rfloor))$
$=\frac{2}{n^{2}}\sum_{i=1}^{p}\frac{(n-\lfloor nt_{1}\rfloor)\lfloor nt_{2}\rfloor}{(n-\lfloor nt_{1}\rfloor-i+1)(\lfloor nt_{2}\rfloor-i+1)}\left\{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor-S_{i-1,1}-S_{i-1,2}+S_{i-1,3}\right\}+o_{\mathbb{P}}(1)$
$=\tau_{0,n}+\tau_{3,2,n}+o_{\mathbb{P}}(1),$ (A.19)
where
$\tau_{0,n}$ $=\frac{2}{n^{2}}\sum_{i=1}^{p}\frac{(n-\lfloor nt_{1}\rfloor)\lfloor nt_{2}\rfloor}{(n-\lfloor nt_{1}\rfloor-i+1)(\lfloor nt_{2}\rfloor-i+1)}\left\{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor-S_{i-1,1}-S_{i-1,2}+S_{i-1,3,1}\right\},$
$\tau_{3,2,n}$ $=\frac{2}{n^{2}}\sum_{i=1}^{p}\frac{(n-\lfloor nt_{1}\rfloor)\lfloor nt_{2}\rfloor}{(n-\lfloor nt_{1}\rfloor-i+1)(\lfloor nt_{2}\rfloor-i+1)}S_{i-1,3,2}.$
To simplify the first term $\tau_{0,n}$, we first note that using (A.12), (A.13), (A.15)
$\frac{1}{n}\Big\{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor-S_{i-1,1}-S_{i-1,2}+S_{i-1,3,1}\Big\}$
$=\frac{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor}{n}\left(1+\frac{(i-1)^{2}}{\lfloor nt_{2}\rfloor(n-\lfloor nt_{1}\rfloor)}-\frac{i-1}{n-\lfloor nt_{1}\rfloor}-\frac{i-1}{\lfloor nt_{2}\rfloor}\right)+o_{\mathbb{P}}(1)$
$=\frac{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor}{n}\frac{(\lfloor nt_{2}\rfloor-i+1)(n-\lfloor nt_{1}\rfloor-i+1)}{\lfloor nt_{2}\rfloor(n-\lfloor nt_{1}\rfloor)}+o_{\mathbb{P}}(1).$
This implies
$\tau_{0,n}$ $=\frac{2(\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor)p}{n^{2}}+o_{\mathbb{P}}(1)=2y(t_{2}-t_{1})+o_{\mathbb{P}}(1).$ (A.20)
Using (A.18), we get for the second term
$\tau_{3,2,n}$ $=\frac{2}{n}\sum_{i=1}^{p}\left(\frac{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor}{n}\right)^{2}\frac{(i-1)n}{\left(n-\lfloor nt_{1}\rfloor\right)\lfloor nt_{2}\rfloor-\left(\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor\right)(i-1)}+o_{\mathbb{P}}(1)$
$=\frac{2(\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor)^{2}}{n^{2}}\sum_{i=1}^{p}\frac{(i-1)}{\left(n-\lfloor nt_{1}\rfloor\right)\lfloor nt_{2}\rfloor-\left(\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor\right)(i-1)}+o_{\mathbb{P}}(1)$
$=\frac{2(\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor)^{2}}{n^{2}}\frac{1}{p}\sum_{i=1}^{p}\frac{(i-1)p}{\left(n-\lfloor nt_{1}\rfloor\right)\lfloor nt_{2}\rfloor-\left(\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor\right)(i-1)}+o_{\mathbb{P}}(1)$
$=2(t_{2}-t_{1})^{2}\int_{0}^{1}\frac{yx}{y^{-1}(1-t_{1})t_{2}-(t_{2}-t_{1})x}dx+o_{\mathbb{P}}(1)$
$=2(t_{2}-t_{1})^{2}\int_{0}^{1}\frac{y^{2}x}{(1-t_{1})t_{2}-(t_{2}-t_{1})yx}dx+o_{\mathbb{P}}(1)$
$=2(t_{2}-t_{1})^{2}\int_{0}^{y}\frac{x}{(1-t_{1})t_{2}-(t_{2}-t_{1})x}dx+o_{\mathbb{P}}(1)$
$=2(t_{2}-t_{1})^{2}\Big[-\frac{(1-t_{1})t_{2}\log\left((1-t_{1})t_{2}-(t_{2}-t_{1})x\right)}{(t_{2}-t_{1})^{2}}-\frac{x}{t_{2}-t_{1}}\Big]_{x=0}^{x=y}+o_{\mathbb{P}}(1)$
$=2(t_{2}-t_{1})\Big\{\frac{(1-t_{1})t_{2}\log\left((1-t_{1})t_{2}\right)}{t_{2}-t_{1}}-\frac{(1-t_{1})t_{2}\log\left((1-t_{1})t_{2}-(t_{2}-t_{1})y\right)}{t_{2}-t_{1}}-y\Big\}+o_{\mathbb{P}}(1)$
$=2(t_{2}-t_{1})\Bigg\{-\frac{(1-t_{1})t_{2}}{t_{2}-t_{1}}\log\left(1-\frac{(t_{2}-t_{1})y}{(1-t_{1})t_{2}}\right)-y\Bigg\}+o_{\mathbb{P}}(1)$
$=-2(1-t_{1})t_{2}\log\left(1-\frac{(t_{2}-t_{1})y}{(1-t_{1})t_{2}}\right)-2y(t_{2}-t_{1})+o_{\mathbb{P}}(1).$ (A.21)
Combining the results for $\tau_{0,n}$ in (A.20) and $\tau_{3,2,n}$ in (A.21) and using (A.19), we get
$\sigma^{2}(\lfloor nt_{1}\rfloor+1,n,1,\lfloor nt_{2}\rfloor))=-2(1-t_{1})t_{2}\log\left(1-\frac{(t_{2}-t_{1})y}{(1-t_{1})t_{2}}\right)+o_{\mathbb{P}}(1),$
which concludes the proof. ∎
Lemma 7.
For $t_{2}>t_{1},$ we have
$\frac{1}{n}\operatorname{tr}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{-1}\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{-1}\right)$
$=\frac{in\left(n-\lfloor nt_{1}\rfloor\right)\lfloor nt_{2}\rfloor}{\left(n-\lfloor nt_{1}\rfloor-i\right)\left(\lfloor nt_{2}\rfloor-i\right)\left\{\left(n-\lfloor nt_{1}\rfloor\right)\lfloor nt_{2}\rfloor-\left(\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor\right)i\right\}}+o_{\mathbb{P}}(1).$
Proof of Lemma 7.
To compute the trace, we use the general strategy of Dörnemann (2022); Dörnemann and Paul (2024). Note that, however, their results do not apply to our situation. Indeed, the terms of interest admit subtle differences and needs to be studied carefully. Similarly to (Dörnemann, 2022, (6.25)), we have the following decomposition for $\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}$,
$\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{-1}=\frac{1}{\frac{n-\lfloor nt_{1}\rfloor}{n}b_{(\lfloor nt_{1}\rfloor+1):n}}\mathbf{I}+b_{(\lfloor nt_{1}\rfloor+1):n}\mathbf{A}+\mathbf{B}+\mathbf{C},$
where
$\mathbf{A}$ $=-\frac{1}{\frac{n-\lfloor nt_{1}\rfloor}{n}b_{i,(\lfloor nt_{1}\rfloor+1):n}}\sum\limits_{\begin{subarray}{c}k=\lfloor nt_{1}\rfloor+1\end{subarray}}^{n}\left(n^{-1}\mathbf{x}_{i,k}\mathbf{x}_{i,k}^{\top}-n^{-1}\mathbf{I}\right)\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{(-k)}\right)^{-1},$
$\mathbf{B}$ $=-\frac{1}{\frac{n-\lfloor nt_{1}\rfloor}{n}b_{i,(\lfloor nt_{1}\rfloor+1):n}}\sum\limits_{\begin{subarray}{c}i=\lfloor nt_{1}\rfloor+1\end{subarray}}^{n}\left(\beta_{i,(\lfloor nt_{1}\rfloor+1):n}^{(-k)}-b_{i,(\lfloor nt_{1}\rfloor+1):n}\right)n^{-1}\mathbf{x}_{i,k}\mathbf{x}_{i,k}^{\top}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{(-k)}\right)^{-1},$
$\mathbf{C}$ $=n^{-1}\sum_{k=\lfloor nt_{1}\rfloor+1}^{n}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{-1}-\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{(-k)}\right)^{-1}\right).$
A similar decomposition can be derived for $\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{-1}$. In the following, we apply this decomposition to $(1/n)\operatorname{tr}\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{-1}\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{-1}$ and to identify the contributing terms. Similarly to the arguments given in Section 6.3.2 (Step 2.1) in Dörnemann (2022), we see that terms involving $\mathbf{B}_{s}$ and $\mathbf{C}_{s}$ are asymptotically negligible, among others. Applying the representation (B.12) in Dörnemann and Paul (2024) to our setting and using (A.10), we get
$n^{-1}\operatorname{tr}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{-1}\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{-1}\right)$ $=A_{t_{1},t_{2}}+\frac{i}{n}\frac{1}{\frac{n-\lfloor nt_{1}\rfloor}{n}b_{i,(\lfloor nt_{1}\rfloor+1):n}\frac{\lfloor nt_{2}\rfloor}{n}b_{i,1:\lfloor nt_{2}\rfloor}}+o_{\mathbb{P}}(1)$
$=A_{t_{1},t_{2}}+\frac{in}{\left(n-\lfloor nt_{1}\rfloor-i\right)\left(\lfloor nt_{2}\rfloor-i\right)}+o_{\mathbb{P}}(1),$
where
$b_{i,(\lfloor nt_{1}\rfloor+1):n}$ $=\frac{1}{1+n^{-1}\mathbb{E}\left[\operatorname{tr}\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{-1}\right]},\quad b_{i,1:\lfloor nt_{2}\rfloor}=\frac{1}{1+n^{-1}\mathbb{E}\left[\operatorname{tr}\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{-1}\right]},$
$A_{t_{1},t_{2}}$ $=\frac{1}{n^{3}}\frac{1}{\frac{n-\lfloor nt_{1}\rfloor}{n}}\sum_{k=\lfloor nt_{1}\rfloor+1}^{\lfloor nt_{2}\rfloor}\beta_{i,1:\lfloor nt_{2}\rfloor}^{(-k)}\mathbf{x}_{i,k}^{\top}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{(-k)}\right)^{-1}\left(\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{(-k)}\right)^{-1}\mathbf{x}_{i,k}\mathbf{x}_{i,k}^{\top}\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{(-k)}\mathbf{x}_{i,k}$
$=\frac{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor}{n^{2}\left(n-\lfloor nt_{1}\rfloor\right)}\beta_{i,1:\lfloor nt_{2}\rfloor}\operatorname{tr}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{-1}\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{-1}\right)\operatorname{tr}\left(\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{-1}\right)+o_{\mathbb{P}}(1)$
$=\frac{\left(\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor\right)i}{\left(n-\lfloor nt_{1}\rfloor\right)\lfloor nt_{2}\rfloor}\frac{1}{n}\operatorname{tr}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{-1}\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{-1}\right)+o_{\mathbb{P}}(1).$
This implies
$n^{-1}\operatorname{tr}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{-1}\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{-1}\right)=\frac{\frac{in}{\left(n-\lfloor nt_{1}\rfloor-i\right)\left(\lfloor nt_{2}\rfloor-i\right)}}{1-\frac{\left(\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor\right)i}{\left(n-\lfloor nt_{1}\rfloor\right)\lfloor nt_{2}\rfloor}}+o_{\mathbb{P}}(1)$
$=\frac{in\left(n-\lfloor nt_{1}\rfloor\right)\lfloor nt_{2}\rfloor}{\left(n-\lfloor nt_{1}\rfloor-i\right)\left(\lfloor nt_{2}\rfloor-i\right)\left\{\left(n-\lfloor nt_{1}\rfloor\right)\lfloor nt_{2}\rfloor-\left(\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor\right)i\right\}}+o_{\mathbb{P}}(1).$
In the following, we prove the approximation for $\sigma_{2}$ appearing in Lemma 3.
Proof of Lemma 3.
By definition of $\sigma_{2}^{2}$, it suffices to show that
$\sum_{i=1}^{p}\frac{\operatorname{tr}\left(\mathbf{P}(i-1;j_{1}:k_{1})\odot\mathbf{P}^{j_{1}:k_{1}}(i-1;j_{2}:k_{2})\right)}{(k_{1}-j_{1}-i+1)(k_{2}-j_{2}-i+1)}=\frac{p}{k_{2}-j_{2}+1}+o_{\mathbb{P}}(1)$ (A.22)
and
$\sum_{i=1}^{p}\frac{\operatorname{tr}\left(\mathbf{P}^{(\lfloor nt_{1}\rfloor+1):\lfloor nt_{2}\rfloor}(i-1;(\lfloor nt_{1}\rfloor+1):n)\odot\mathbf{P}^{(\lfloor nt_{1}\rfloor+1):\lfloor nt_{2}\rfloor}(i-1;1:\lfloor nt_{2}\rfloor)\right)}{(\lfloor nt_{2}\rfloor-i+1)(n-\lfloor nt_{1}\rfloor-i+1)}$
$\quad=\frac{p(\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor)}{\lfloor nt_{2}\rfloor(n-\lfloor nt_{1}\rfloor)}+o_{\mathbb{P}}(1).$ (A.23)
We begin with a proof of (A.23). Note that one can show similarly to (A.12)
$\frac{1}{n-\lfloor nt_{1}\rfloor}\sum_{k=\lfloor nt_{1}\rfloor+1}^{\lfloor nt_{2}\rfloor}\left(\frac{\mathbf{P}(i-1;1:\lfloor nt_{2}\rfloor)}{\lfloor nt_{2}\rfloor-i+1}\right)_{kk}$ $=\frac{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor}{\lfloor nt_{2}\rfloor(n-\lfloor nt_{1}\rfloor)}+o_{\mathbb{P}}(1),$
$\frac{1}{n-\lfloor nt_{1}\rfloor}\sum_{k=\lfloor nt_{1}\rfloor+1}^{\lfloor nt_{2}\rfloor}\left(\frac{\mathbf{P}(i-1;1:\lfloor nt_{2}\rfloor)}{\lfloor nt_{2}\rfloor-i+1}\right)_{kk}$ $=\frac{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor}{\lfloor nt_{2}\rfloor(n-\lfloor nt_{1}\rfloor)}+o_{\mathbb{P}}(1).$
This gives
$\operatorname{tr}\Bigg\{\left(\frac{\mathbf{P}^{(\lfloor nt_{1}\rfloor+1):\lfloor nt_{2}\rfloor}(i-1;(\lfloor nt_{1}\rfloor+1):n)}{n-\lfloor nt_{1}\rfloor-i+1}-\frac{1}{n-\lfloor nt_{1}\rfloor}\mathbf{I}_{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor}\right)$
$\quad\quad\odot\left(\frac{\mathbf{P}^{(\lfloor nt_{1}\rfloor+1):\lfloor nt_{2}\rfloor}(i-1;1:\lfloor nt_{2}\rfloor)}{\lfloor nt_{2}\rfloor-i+1}-\frac{1}{\lfloor nt_{2}\rfloor}\mathbf{I}_{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor}\right)\Bigg\}$
$=\frac{\operatorname{tr}\left(\mathbf{P}^{(\lfloor nt_{1}\rfloor+1):\lfloor nt_{2}\rfloor}(i-1;(\lfloor nt_{1}\rfloor+1):n)\odot\mathbf{P}^{(\lfloor nt_{1}\rfloor+1):\lfloor nt_{2}\rfloor}(i-1;1:\lfloor nt_{2}\rfloor)\right)}{(\lfloor nt_{2}\rfloor-i+1)(n-\lfloor nt_{1}\rfloor-i+1)}$
$\quad-\frac{1}{n-\lfloor nt_{1}\rfloor}\sum_{k=\lfloor nt_{1}\rfloor+1}^{\lfloor nt_{2}\rfloor}\left(\frac{\mathbf{P}(i-1;1:\lfloor nt_{2}\rfloor)}{\lfloor nt_{2}\rfloor-i+1}\right)_{kk}$
$\quad-\frac{1}{\lfloor nt_{2}\rfloor}\sum_{k=\lfloor nt_{1}\rfloor+1}^{\lfloor nt_{2}\rfloor}\left(\frac{\mathbf{P}(i-1;(\lfloor nt_{1}\rfloor+1):n)}{n-\lfloor nt_{1}\rfloor-i+1}\right)_{kk}+\frac{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor}{\lfloor nt_{2}\rfloor(n-\lfloor nt_{1}\rfloor)}$
$=\frac{\operatorname{tr}\left(\mathbf{P}^{(\lfloor nt_{1}\rfloor+1):\lfloor nt_{2}\rfloor}(i-1;(\lfloor nt_{1}\rfloor+1):n)\odot\mathbf{P}^{(\lfloor nt_{1}\rfloor+1):\lfloor nt_{2}\rfloor}(i-1;1:\lfloor nt_{2}\rfloor)\right)}{(\lfloor nt_{2}\rfloor-i+1)(n-\lfloor nt_{1}\rfloor-i+1)}-\frac{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor}{\lfloor nt_{2}\rfloor(n-\lfloor nt_{1}\rfloor)}$
$\quad+o_{\mathbb{P}}(1).$
Using the same arguments as in the proofs of Lemma 4 and 5 in Dörnemann (2023), we conclude that
$\sum_{i=1}^{p}\left\{\frac{\operatorname{tr}\left(\mathbf{P}^{(\lfloor nt_{1}\rfloor+1):\lfloor nt_{2}\rfloor}(i-1;(\lfloor nt_{1}\rfloor+1):n)\odot\mathbf{P}^{(\lfloor nt_{1}\rfloor+1):\lfloor nt_{2}\rfloor}(i-1;1:\lfloor nt_{2}\rfloor)\right)}{(\lfloor nt_{2}\rfloor-i+1)(n-\lfloor nt_{1}\rfloor-i+1)}-\frac{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor}{\lfloor nt_{2}\rfloor(n-\lfloor nt_{1}\rfloor)}\right\}$
$=\sum_{i=1}^{p}\operatorname{tr}\Bigg\{\left(\frac{\mathbf{P}^{(\lfloor nt_{1}\rfloor+1):\lfloor nt_{2}\rfloor}(i-1;(\lfloor nt_{1}\rfloor+1):n)}{n-\lfloor nt_{1}\rfloor-i+1}-\frac{1}{n-\lfloor nt_{1}\rfloor}\mathbf{I}_{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor}\right)$
$\quad\quad\quad\quad\odot\left(\frac{\mathbf{P}^{(\lfloor nt_{1}\rfloor+1):\lfloor nt_{2}\rfloor}(i-1;1:\lfloor nt_{2}\rfloor)}{\lfloor nt_{2}\rfloor-i+1}-\frac{1}{\lfloor nt_{2}\rfloor}\mathbf{I}_{\lfloor nt_{2}\rfloor-\lfloor nt_{1}\rfloor}\right)\Bigg\}+o_{\mathbb{P}}(1)$
$=o_{\mathbb{P}}(1),$
which implies (A.23). The assertion (A.22) can be shown very similarly and is omitted for the sake of brevity. ∎
We are now in the position to prove the following auxiliary result on the quadratic term given previously in Lemma 1.
Proof of Lemma 1.
Define
$A_{i,t}=\frac{\lfloor nt\rfloor}{n}X_{i,1:\lfloor nt\rfloor}^{2}+\frac{n-\lfloor nt\rfloor}{n}X_{i,(\lfloor nt\rfloor+1):n}^{2}-X_{i}^{2},\quad 1\leq i\leq p,t\in[t_{0},1-t_{0}].$ (A.24)
To begin with, we show that
$\sum_{i=1}^{p}\mathbb{E}[A_{i,t}|\mathcal{A}_{i-1}]-\breve{\sigma}^{2}_{n,t}=o_{\mathbb{P}}(1).$ (A.25)
Recalling (4.27), (4.28) and (4.29), we see that
$\sum\limits_{i=1}^{p}\mathbb{E}[A_{i,t}|\mathcal{A}_{i-1}]$ $=\frac{n}{\lfloor nt\rfloor}\sigma^{2}(1,\lfloor nt\rfloor,1,\lfloor nt\rfloor)+\frac{n}{n-\lfloor nt\rfloor}\sigma^{2}(\lfloor nt\rfloor+1,n,\lfloor nt\rfloor+1,n)-\sigma^{2}(1,n,1,n)$
$=2\log\left(1-\frac{p}{n}\right)-2\frac{\lfloor nt\rfloor}{n}\log\left(1-\frac{p}{\lfloor nt\rfloor}\right)-2\frac{n-\lfloor nt\rfloor}{n}\log\left(1-\frac{p}{n-\lfloor nt\rfloor}\right)$
$\quad+\frac{(\mathbb{E}[x_{11}^{4}]-3)p}{n}+o_{\mathbb{P}}(1)$
$=\breve{\sigma}_{n,t}^{2}+o_{\mathbb{P}}(1),$ (A.26)
where we used Proposition 1. This implies assertion (A.25). Thus, it remains to show that
$\sum_{i=1}^{p}\left(A_{i,t}-\mathbb{E}[A_{i,t}|\mathcal{A}_{i-1}]\right)=o_{\mathbb{P}}(1),$
which follows from (A.2) and (A.47) given later. ∎
A.4 Proof of Proposition 1
Define
$\tau_{n}$ $=\|\boldsymbol{\Sigma}_{n}\|_{F}^{2},$
$\nu_{n}$ $=\operatorname{Var}\left(\|\mathbf{y}_{1}-\mathbb{E}[\mathbf{y}_{1}]\|_{2}^{2}\right),$
$\omega_{n}$ $=\sum_{j=1}^{p}\Sigma_{jj}^{2}.$
Then, (2.6) can be written as
$\kappa_{n}=3+\frac{\nu_{n}-2\tau_{n}}{\omega_{n}}.$
[…]
Lemma 8.
Suppose that assumptions (A-1) and (A-2) are satisfied, and that $H_{0}$ holds true. Then, it holds that
$\frac{\hat{\tau}_{n}}{\tau_{n}}\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}1.$ (a)
$\frac{1}{\omega_{n}}\mathbb{E}\left|\hat{\omega}_{n}-\omega_{n}\right|\to 0,$ (b)
$\frac{\hat{\nu}_{n}}{\nu_{n}}\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}1.$ (c)
Proof of Lemma 8.
For the proof of (a), we refer to (Bai and Saranadasa, 1996, Section A.3).
To prove (b), we define
$\overline{y}_{j\cdot}$ $=\frac{1}{n}\sum_{i^{\prime}=1}^{n}y_{ji^{\prime}},$
$\hat{\sigma}_{j}^{2}$ $=\frac{1}{n}\sum_{i=1}^{n}\left(y_{ji}-\overline{y}_{j\cdot}\right)^{2},$
$\hat{\sigma}_{j,1}^{2}$ $=\frac{1}{n}\sum_{i=1}^{n}\left(y_{ji}-\mathbb{E}[y_{ji}]\right)^{2},$
$\hat{\sigma}_{j,2}^{2}$ $=\left(\mathbb{E}[y_{j1}]-\overline{y}_{j\cdot}\right)^{2},\quad 1\leq j\leq p.$
Then, we have for $\hat{\omega}_{n}$ that
$\hat{\omega}_{n}=\sum_{j=1}^{p}\left(\hat{\sigma}_{j}^{2}\right)^{2}\lesssim\sum_{j=1}^{p}\left(\hat{\sigma}_{j,1}^{2}\right)^{2}+\sum_{j=1}^{p}\left(\hat{\sigma}_{j,2}^{2}\right)^{2}=:\hat{\omega}_{n,1}+\hat{\omega}_{n,2}.$
Then, it follows from Lemma S.2 in Lopes et al. (2019) that
$\frac{1}{\omega_{n}}\mathbb{E}\left|\hat{\omega}_{n,1}-\omega_{n}\right|\to 0.$ (A.27)
We continue with studying the second term $\hat{\omega}_{n,2}.$ Without loss of generality, we may assume that $\mathbb{E}[y_{j1}]=0$ for all $1\leq j\leq p$, and we use the notation $(U_{kl})_{1\leq k,l\leq p}=\boldsymbol{\Sigma}^{1/2}.$ As a preparation, we note that $\mathbb{E}[y_{j1}^{2}]=\Sigma_{jj}\leq\|\boldsymbol{\Sigma}\|\lesssim 1$. Moreover, note that $\max_{1\leq k,l\leq p}|U_{kl}|\lesssim 1$, where $U_{kl}$ denote the entries of $\boldsymbol{\Sigma}^{1/2}$. Then, one can also verify by a direct calculation $\mathbb{E}[y_{j1}^{4}]\lesssim 1$. These considerations imply
$\mathbb{E}[\overline{y}_{j\cdot}^{2}]$ $=\frac{1}{n^{2}}\sum_{i=1}^{n}\mathbb{E}[y_{ji}^{2}]=\frac{1}{n}\Sigma_{jj}\lesssim\frac{1}{n},$ (A.28)
$\mathbb{E}[\overline{y}_{j\cdot}^{4}]$ $\lesssim\frac{1}{n^{3}}\mathbb{E}[y_{j1}^{4}]+\frac{1}{n^{2}}\left(\mathbb{E}[y_{j1}^{2}]\right)^{2}\lesssim\frac{1}{n^{2}}.$ (A.29)
Then, we obtain
$\mathbb{E}\left(\hat{\sigma}_{j,2}^{2}\right)^{2}$ $=\mathbb{E}\left(\overline{y}_{j\cdot}\right)^{4}\lesssim\frac{1}{n^{2}}.$
As $\omega_{n}\gtrsim 1$, we conclude that
$\frac{1}{\omega_{n}}\mathbb{E}\left|\hat{\omega}_{n,2}\right|=\frac{1}{\omega_{n}}\sum_{j=1}^{p}\mathbb{E}\left(\hat{\sigma}_{j,2}^{2}\right)^{2}=o(1).$ (A.30)
Then, assertion (b) follows from (A.27) and (A.30).
For a proof of part (c) we note Lemma S.3 in Lopes et al. (2019) implies
$\frac{\breve{\nu}_{n}}{\nu_{n}}\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}1,$ (A.31)
where
$\breve{\nu}_{n}=\frac{1}{n-1}\sum_{i=1}^{n}\left(\|\mathbf{y}_{i}-\mathbb{E}[\mathbf{y}_{i}]\|_{2}^{2}-\frac{1}{n}\sum_{i=1}^{n}\|\mathbf{y}_{i}-\mathbb{E}[\mathbf{y}_{i}]\|_{2}^{2}\right)^{2}.$
Then, (c) follows from (A.31) and
$\mathbb{E}\left|\frac{\breve{\nu}_{n}-\hat{\nu}_{n}}{\nu_{n}}\right|=o(1).$ (A.32)
In the following, we will verify (A.32) assuming w.l.o.g. that $\mathbb{E}[x_{11}]=0.$ We define
$\breve{\nu}_{n,1}^{1/2}$ $=\|\mathbf{y}_{1}\|_{2}^{2}-\frac{1}{n}\sum_{i=1}^{n}\|\mathbf{y}_{i}\|_{2}^{2},$
$\hat{\nu}_{n,1}^{1/2}$ $=\|\mathbf{y}_{1}-\overline{\mathbf{y}}\|_{2}^{2}-\frac{1}{n}\sum_{i=1}^{n}\|\mathbf{y}_{i}-\overline{\mathbf{y}}\|_{2}^{2}.$
Step 1 Let $\overline{y}_{1\cdot},\ldots,\overline{y}_{p\cdot}$ denote the components of the $p$-dimensional vector $\overline{\mathbf{y}}.$ Then, a direct computation gives
$\mathbb{E}(\breve{\nu}_{n,1}^{1/2}-\hat{\nu}_{n,1}^{1/2})^{2}$ $=\mathbb{E}\Big[2\sum_{j=1}^{p}\overline{y}_{j\cdot}\big(y_{j1}-\overline{y}_{j\cdot}\big)\Big]^{2}=\sum_{j=1}^{p}\mathbb{E}[T_{1,j}]+\sum_{\begin{subarray}{c}j,k=1,\\ j\neq k\end{subarray}}^{p}\mathbb{E}[T_{2,j,k}],$ (A.33)
where for $1\leq j\neq k\leq p$
$T_{1,j}$ $=\overline{y}_{j\cdot}^{2}\left(y_{j1}-\overline{y}_{j\cdot}\right)^{2}$
$T_{2,j,k}$ $=\overline{y}_{j\cdot}\overline{y}_{k\cdot}\left(y_{j1}-\overline{y}_{j\cdot}\right)\left(y_{k1}-\overline{y}_{k\cdot}\right)$
In the following, we use the notation
$\overline{y}_{j,-1}=\frac{1}{n}\sum_{i=2}^{n}y_{ji}=\overline{y}_{j\cdot}-\frac{1}{n}y_{j1},$ (A.34)
which is independent of $y_{j1},~1\leq j\leq p.$ Subsequently, we analyze $T_{1,j}$ and $T_{2,j,k}.$ For the mean of the first term, we use (A.28) and (A.29) to get
$\mathbb{E}[T_{1,j}]$ $=\mathbb{E}\left[\overline{y}_{j\cdot}^{2}\left(y_{j1}-\overline{y}_{j\cdot}\right)^{2}\right]\lesssim\mathbb{E}[\overline{y}_{j\cdot}^{2}y_{j1}^{2}]+\mathbb{E}[\overline{y}_{j\cdot}^{4}]\lesssim\mathbb{E}\left[\left(\overline{y}_{j,-1}+n^{-1}y_{j1}\right)^{2}y_{j1}^{2}\right]+n^{-1}$
$\lesssim\mathbb{E}\left[\overline{y}_{j,-1}^{2}y_{j1}^{2}\right]+n^{-2}\mathbb{E}\left[y_{j1}^{4}\right]+n^{-1}\lesssim\mathbb{E}\left[\overline{y}_{j,-1}^{2}\right]\mathbb{E}\left[y_{j1}^{2}\right]+n^{-1}\lesssim n^{-1}.$ (A.35)
For the mean of the second term, we expand the brackets and get
$\mathbb{E}[T_{2,j,k}]$ $=\mathbb{E}[T_{2,1,j,k}]-\mathbb{E}[T_{2,2,j,k}]-\mathbb{E}[T_{2,3,j,k}]+\mathbb{E}[T_{2,4,j,k}],$ (A.36)
where
$T_{2,1,j,k}$ $=\overline{y}_{j\cdot}\overline{y}_{k\cdot}y_{j1}y_{k1},$
$T_{2,2,j,k}$ $=\overline{y}_{j\cdot}\overline{y}_{k\cdot}^{2}y_{j1},$
$T_{2,3,j,k}$ $=\overline{y}_{j\cdot}^{2}\overline{y}_{k\cdot}y_{k1},$
$T_{2,4,j,k}$ $=\overline{y}_{j\cdot}^{2}\overline{y}_{k\cdot}^{2}.$
Using (A.34), we get
$\left|\mathbb{E}[T_{2,2,j,k}]\right|$ $\leq\frac{1}{n}\mathbb{E}[\overline{y}_{k\cdot}^{2}y_{j1}^{2}]+\left|\mathbb{E}[\overline{y}_{j,-1}\overline{y}_{k\cdot}^{2}y_{j1}]\right|\lesssim\frac{1}{n^{2}}+\left|\mathbb{E}[\overline{y}_{j,-1}\overline{y}_{k\cdot}^{2}y_{j1}]\right|$
$\leq\frac{1}{n^{2}}+\left|\mathbb{E}[\overline{y}_{j,-1}\overline{y}_{k,-1}^{2}y_{j1}]\right|+\frac{1}{n^{2}}\left|\mathbb{E}[\overline{y}_{j,-1}y_{k1}^{2}y_{j1}]\right|+\frac{2}{n}\left|\mathbb{E}[\overline{y}_{j,-1}\overline{y}_{k,-1}y_{j1}y_{k1}]\right|$
$\leq\frac{1}{n^{2}}+\frac{2}{n}\left|\mathbb{E}[\overline{y}_{j,-1}\overline{y}_{k,-1}]\mathbb{E}[y_{j1}y_{k1}]\right|$
$\lesssim\frac{1}{n^{2}},$ (A.37)
where we used that (as a consequence of (A.28) and (A.29))
$\mathbb{E}[\overline{y}_{k\cdot}^{2}y_{j1}^{2}]$ $\leq\left(\mathbb{E}[\overline{y}_{k\cdot}^{4}]\mathbb{E}[y_{j1}^{4}]\right)^{1/2}\lesssim\frac{1}{n},$
$\mathbb{E}[\overline{y}_{j,-1}\overline{y}_{k,-1}^{2}y_{j1}]$ $=\mathbb{E}[\overline{y}_{j,-1}\overline{y}_{k,-1}^{2}]\mathbb{E}[y_{j1}]=0,$
$\mathbb{E}[\overline{y}_{j,-1}y_{k1}^{2}y_{j1}]$ $=\mathbb{E}[\overline{y}_{j,-1}]\mathbb{E}[y_{k1}^{2}y_{j1}]=0,$
$\left|\mathbb{E}[\overline{y}_{j,-1}\overline{y}_{k,-1}]\mathbb{E}[y_{j1}y_{k1}]\right|$ $=\left|\Sigma_{kj}\mathbb{E}[\overline{y}_{j,-1}\overline{y}_{k,-1}]\right|\lesssim\left(\mathbb{E}[\overline{y}_{j,-1}^{2}]\mathbb{E}[\overline{y}_{k,-1}^{2}]\right)^{1/2}\lesssim\frac{1}{n}.$
Similarly to the considerations for $T_{2,2,j,k}$, we get
$\left|\mathbb{E}[T_{2,3,j,k}]\right|\lesssim\frac{1}{n^{2}}.$ (A.38)
By an application of Hölder’s inequality and (A.29), we get
$\mathbb{E}[T_{2,4,j,k}]\lesssim\frac{1}{n^{2}}.$ (A.39)
It is left to analyze the mean of the term $T_{2,1,j,k}$. Using (A.34) and the fact $\mathbb{E}[\overline{y}_{j,-1}]=0$ for all $1\leq j\leq p$, we obtain
$\mathbb{E}[T_{2,1,j,k}]=\mathbb{E}\left[\left(\overline{y}_{j,-1}+\frac{1}{n}y_{j1}\right)\left(\overline{y}_{k,-1}+\frac{1}{n}y_{k1}\right)y_{j1}y_{k1}\right]$
$=\mathbb{E}[\overline{y}_{j,-1}\overline{y}_{k,-1}]\mathbb{E}[y_{j1}y_{k1}]+\frac{1}{n}\mathbb{E}\left[\overline{y}_{j,-1}\right]\mathbb{E}\left[y_{j1}y_{k1}^{2}\right]+\frac{1}{n}\mathbb{E}\left[\overline{y}_{k,-1}\right]\mathbb{E}\left[y_{j1}^{2}y_{k1}\right]+\frac{1}{n^{2}}\mathbb{E}\left[y_{j1}^{2}y_{k1}^{2}\right]$
$=\mathbb{E}[\overline{y}_{j,-1}\overline{y}_{k,-1}]\Sigma_{j,k}+\frac{1}{n^{2}}\mathbb{E}\left[y_{j1}^{2}y_{k1}^{2}\right].$
Note that
$\mathbb{E}[\overline{y}_{j,-1}\overline{y}_{k,-1}]=\frac{1}{n^{2}}\sum_{i,i^{\prime}=2}^{n}\mathbb{E}[y_{ji}y_{ki^{\prime}}]=\frac{1}{n^{2}}\sum_{i=2}^{n}\mathbb{E}[y_{ji}y_{ki}]=\frac{n-1}{n^{2}}\Sigma_{kj}$
This implies
$\left|\mathbb{E}[T_{2,1,j,k}]\right|\lesssim|\Sigma_{jk}|\left|\mathbb{E}[\overline{y}_{j,-1}\overline{y}_{k,-1}]\right|+\frac{1}{n^{2}}\lesssim\frac{1}{n}\Sigma_{kj}^{2}+\frac{1}{n^{2}}.$ (A.40)
Combining (A.36) with the bounds (A.37), (A.38), (A.39), (A.40), we obtain
$\left|\mathbb{E}[T_{2,j,k}]\right|\lesssim\frac{1}{n^{2}}+\frac{1}{n}\Sigma_{jk}^{2}.$ (A.41)
In summary, we obtain using (A.33), (A.35), (A.41) and assumption (A-3)
[…]
Step 2 Note that $\breve{\nu}_{n}$ is unbiased for $\nu_{n}.$ Therefore, we get
$\mathbb{E}\left[\frac{\breve{\nu}_{n,1}}{\nu_{n}}\right]=\frac{n-1}{n\nu_{n}}\mathbb{E}\left[\breve{\nu}_{n}\right]=\frac{n-1}{n}\leq 1.$ (A.43)
From (A.42) and (A.43), we also obtain
$\mathbb{E}\left[\frac{\hat{\nu}_{n,1}}{\nu_{n}}\right]\lesssim\mathbb{E}\left[\frac{\left(\hat{\nu}_{n,1}^{1/2}-\breve{\nu}_{n,1}^{1/2}\right)^{2}}{\nu_{n}}\right]+\mathbb{E}\left[\frac{\breve{\nu}_{n,1}}{\nu_{n}}\right]\lesssim 1.$ (A.44)
Conclusion Using (A.42), (A.43), (A.44) and $\nu_{n}\gtrsim n$ (see p.3 in the supplementary material of Lopes et al., 2019), we obtain
$\mathbb{E}\left|\frac{\breve{\nu}_{n}-\hat{\nu}_{n}}{\nu_{n}}\right|$ $\lesssim\nu_{n}^{-1}\mathbb{E}\left|\breve{\nu}_{n,1}-\hat{\nu}_{n,1}\right|=\nu_{n}^{-1}\mathbb{E}\left|(\breve{\nu}_{n,1}^{1/2}-\hat{\nu}_{n,1}^{1/2})(\breve{\nu}_{n,1}^{1/2}+\hat{\nu}_{n,1}^{1/2})\right|$
$\leq\nu_{n}^{-1}\left(\mathbb{E}(\breve{\nu}_{n,1}^{1/2}-\hat{\nu}_{n,1}^{1/2})^{2}\mathbb{E}(\breve{\nu}_{n,1}^{1/2}+\hat{\nu}_{n,1}^{1/2})^{2}\right)^{1/2}$
$\lesssim\Bigg\{\frac{\mathbb{E}(\breve{\nu}_{n,1}^{1/2}-\hat{\nu}_{n,1}^{1/2})^{2}}{\nu_{n}}\left(\frac{\mathbb{E}\breve{\nu}_{n,1}}{\nu_{n}}+\frac{\mathbb{E}\hat{\nu}_{n,1}}{\nu_{n}}\right)\Bigg\}^{1/2}$
$\lesssim\Bigg\{\frac{\mathbb{E}(\breve{\nu}_{n,1}^{1/2}-\hat{\nu}_{n,1}^{1/2})^{2}}{\nu_{n}}\Bigg\}^{1/2}=o(1),$
which implies (A.32). ∎
A.5 Proofs of Lemma 5 - Lemma 6
Proof of Lemma 5.
W.l.o.g. assume that $\lfloor nt_{1}\rfloor>\lfloor nt_{2}\rfloor.$ To begin with, we write
$D_{i,1}-D_{i,2}$ $=\lfloor nt_{1}\rfloor X_{i,1:\lfloor nt_{1}\rfloor}-\lfloor nt_{2}\rfloor X_{i,1:\lfloor nt_{2}\rfloor}+(n-\lfloor nt_{1}\rfloor)X_{i,(\lfloor nt_{1}\rfloor+1):n}-(n-\lfloor nt_{2}\rfloor)X_{i,(\lfloor nt_{2}\rfloor+1):n}$
$=Z_{1}+Z_{2},$
where the random variable $Z_{1}$ and $Z_{2}$ are defined by $Z_{1}=Z_{1,1}+Z_{1,2}$, $Z_{2}=Z_{2,1}+Z_{2,2}$ and
$nZ_{1,1}$ $=\left(\lfloor nt_{1}\rfloor-\lfloor nt_{2}\rfloor\right)\sum_{i=1}^{p}X_{i,1:\lfloor nt_{1}\rfloor},\quad nZ_{1,2}=(-\lfloor nt_{1}\rfloor+\lfloor nt_{2}\rfloor)\sum_{i=1}^{p}X_{i,(\lfloor nt_{1}\rfloor+1):n},$
$nZ_{2,1}$ $=\lfloor nt_{2}\rfloor\sum_{i=1}^{p}\left(X_{i,1:\lfloor nt_{1}\rfloor}-X_{i,1:\lfloor nt_{2}\rfloor}\right),\quad nZ_{2,2}=(n-\lfloor nt_{2}\rfloor)\sum_{i=1}^{p}\left(X_{i,(\lfloor nt_{1}\rfloor+1):n}-X_{i,(\lfloor nt_{2}\rfloor+1):n}\right).$
For reasons of symmetry, we restrict ourselves to a proof of the estimates
$\mathbb{E}[Z_{1,1}^{2}]$ $\lesssim\left|\frac{\lfloor nt_{1}\rfloor-\lfloor nt_{2}\rfloor}{n}\right|^{1+d},\quad\mathbb{E}[|Z_{2,1}|^{2+\delta/2}]\lesssim\left|\frac{\lfloor nt_{1}\rfloor-\lfloor nt_{2}\rfloor}{n}\right|^{1+d}.$ (A.45)
Using formuala (9.8.6) in Bai and Silverstein (2010), we get for the second moment of $Z_{1,1}$
$\mathbb{E}[Z_{1,1}^{2}]$ $=\left(\frac{\lfloor nt_{1}\rfloor-\lfloor nt_{2}\rfloor}{n}\right)^{2}\sum_{i=1}^{p}\mathbb{E}[X_{i,1:\lfloor nt_{1}\rfloor}^{2}]\lesssim\left(\frac{\lfloor nt_{1}\rfloor-\lfloor nt_{2}\rfloor}{n}\right)^{2},$
which proves the first assertion in (A.45). For a proof of the second estimate let $\tilde{\mathbf{P}}(i-1;1:\lfloor nt_{2}\rfloor)$ denote a $\lfloor nt_{1}\rfloor\times\lfloor nt_{1}\rfloor$-matrix with entries
$\left(\tilde{\mathbf{P}}(i-1;1:\lfloor nt_{2}\rfloor)\right)_{ij}=\begin{cases}\left(\mathbf{P}(i-1;1:\lfloor nt_{2}\rfloor)\right)_{ij}~&\textnormal{if }1\leq i,j\leq\lfloor nt_{2}\rfloor,\\ 0&\textnormal{else},\end{cases}\quad 1\leq i,j\leq\lfloor nt_{1}\rfloor.$
By Lemma 2.1 in Li (2003) and Lemma B.26 in Bai and Silverstein (2010), we obtain for $Z_{2,1}$
$\mathbb{E}[|Z_{2,1}|^{2+\delta/2}]$ $\lesssim p^{\delta/4}\sum_{i=1}^{p}\mathbb{E}\left|X_{i,1:\lfloor nt_{1}\rfloor}-X_{i,1:\lfloor nt_{2}\rfloor}\right|^{2+\delta/2}$
$=$ $p^{\delta/4}\sum_{i=1}^{p}\mathbb{E}\left|\mathbf{b}_{i,1:\lfloor nt_{1}\rfloor}^{\top}\left(\frac{\mathbf{P}(i-1;1:\lfloor nt_{1}\rfloor)}{\lfloor nt_{1}\rfloor-i+1}-\frac{\mathbf{P}(i-1;1:\lfloor nt_{2}\rfloor)}{\lfloor nt_{2}\rfloor-i+1}\right)\mathbf{b}_{i,1:\lfloor nt_{1}\rfloor}^{\top}\right|^{2+\delta/2}$
$\lesssim\sum_{i=1}^{p}p^{\delta/4}\left\{\operatorname{tr}\left(\frac{\mathbf{P}(i-1;1:\lfloor nt_{1}\rfloor)}{\lfloor nt_{1}\rfloor-i+1}-\frac{\tilde{\mathbf{P}}(i-1;1:\lfloor nt_{2}\rfloor)}{\lfloor nt_{2}\rfloor-i+1}\right)^{2}\right\}^{1+\delta/4}.$ (A.46)
Note that
$\operatorname{tr}\left(\frac{\mathbf{P}(i-1;1:\lfloor nt_{1}\rfloor)}{\lfloor nt_{1}\rfloor-i+1}-\frac{\tilde{\mathbf{P}}(i-1;1:\lfloor nt_{2}\rfloor)}{\lfloor nt_{2}\rfloor-i+1}\right)^{2}$
$=\frac{1}{\lfloor nt_{1}\rfloor-i+1}+\frac{1}{\lfloor nt_{2}\rfloor-i+1}-2\frac{1}{\lfloor nt_{1}\rfloor-i+1}$
$=\frac{\lfloor nt_{1}\rfloor-\lfloor nt_{2}\rfloor}{(\lfloor nt_{1}\rfloor-i+1)(\lfloor nt_{2}\rfloor-i+1)}\lesssim\frac{\lfloor nt_{1}\rfloor-\lfloor nt_{2}\rfloor}{n^{2}}.$
Combining this with (A.46), the second statement in (A.45) follows. ∎
Proof of Lemma 6.
We define
$Q_{n,1,t}$ $=\sum_{i=1}^{p}\mathbb{E}[A_{i,t}|\mathcal{A}_{i-1}]-\breve{\sigma}^{2}_{n,t},\quad Q_{n,2,t}=\sum_{i=1}^{p}\left(A_{i,t}-\mathbb{E}[A_{i,t}|\mathcal{A}_{i-1}]\right),$ (A.47)
where $A_{it}$ is defined in (A.24). Then, the decomposition (A.1) is obviously true. Note that the definition of $\breve{\sigma}_{n,t}^{2}$ in (4.20) implies
$\sup_{\begin{subarray}{c}t\in[t_{0},1-t_{0}],\\ n\in\mathbb{N}\end{subarray}}\breve{\sigma}_{n,t}^{2}\lesssim 1,$
and that
$\sup_{\begin{subarray}{c}t\in[t_{0},1-t_{0}],\\ n\in\mathbb{N}\end{subarray}}\left|\sum_{i=1}^{p}\mathbb{E}[A_{i,t}|\mathcal{A}_{i-1}]\right|\lesssim 1$
almost surely. Thus, we conclude that $(Q_{n,1,t})$ is asymptotically tight in the space $\ell^{\infty}([t_{0},1-t_{0}])$, and it remains to show (A.2). Applying Lemma 2.2 in Li (2003), we obtain
$\mathbb{E}|Q_{2,n,t}|^{2+\delta/4}$ $\lesssim p^{1+\delta/8}\max_{1\leq i\leq p}\mathbb{E}\left[\left|A_{i,t}-\mathbb{E}[A_{i,t}|\mathcal{A}_{i-1}]\right|^{2+\delta/4}\right]$
$\lesssim p^{1+\delta/8}\max_{1\leq i\leq p}\mathbb{E}\left[|X_{i,1:\lfloor nt\rfloor}|^{4+\delta/2}+|X_{i,(\lfloor nt\rfloor+1):n}|^{4+\delta/2}+|X_{i}|^{4+\delta/2}\right],$
and Lemma B.26 in Bai and Silverstein (2010) have
$\mathbb{E}\left[|X_{i,1:\lfloor nt\rfloor}|^{4+\delta/2}\right]\lesssim\frac{\left[\operatorname{tr}\left\{\mathbf{P}(i-1;1:\lfloor nt\rfloor)\right\}^{2}\right]^{2+\delta/4}}{\left(\lfloor nt\rfloor-i+1\right)^{4+\delta/2}}=\frac{1}{\left(\lfloor nt\rfloor-i+1\right)^{2+\delta/4}}\lesssim\frac{1}{n^{2+\delta/4}},$ (A.48)
uniformly with respect to $1\leq i\leq p$ and $t\in[t_{0},1-t_{0}]$. Similarly, one can show that
$\mathbb{E}\left[|X_{i,(\lfloor nt\rfloor+1):n}|^{2+\delta/2}+|X_{i}|^{2+\delta/2}\right]\lesssim\frac{1}{n^{2+\delta/4}},$ (A.49)
uniformly with respect to $1\leq i\leq p$ and $t\in[t_{0},1-t_{0}]$. Finally, (A.48) and (A.49) imply
$\sup_{t\in[t_{0},1-t_{0}]}\mathbb{E}|Q_{2,n,t}|^{2+\delta/4}\lesssim\frac{1}{n^{1+\delta/8}},$
and the assertion of Lemma 6 follows. ∎